### v1

# Variable Neighborhood Searching Rerandomization

Abstract
Rerandomization discards undesired treatment assignments to ensure covariate balance in randomized experiments. However, rerandomization based on acceptance-rejection sampling is computationally inefficient, especially when numerous independent assignments are required to perform randomization-based statistical inference. Existing acceleration methods are suboptimal and are not applicable in structured experiments, including stratified experiments and experiments with clusters. Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to draw balanced assignments in various experiments efficiently. We derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR. Simulation studies and a real data example indicate that our method maintains the appealing statistical properties of rerandomization and can sample thousands of treatment assignments within seconds, even in cases where existing methods require an hour to complete the task.
Keywords causal inference  $\cdot$ clinical trials  $\cdot$ experimental design  $\cdot$ combinatorial optimization  $\cdot$ binary quadratic programming  $\cdot$ variable neighborhood search
1 Introduction
Randomized experiments balance covariates on average and have been the gold standard for drawing causal inferences. However, as pointed out by Morgan and Rubin (2012), covariate imbalances between treatment groups often exist in practice due to the randomness in complete randomization(CR). An intuitive and practical solution to this issue is randomizing repeatedly until a prespecified covariate balance criterion is achieved. Despite the long history and widespread use of rerandomization (Student, 1938; Cox, 1982; Bailey and Rowley, 1987; Maclure et al., 2006; Imai et al., 2008; Bruhn and McKenzie, 2009), its theoretical implications were first studied by Morgan and Rubin (2012) using the Mahalanobis distance only recently. Since then, rerandomization has attracted growing research interests, and its theory underpinnings have been established in various scenarios (Morgan and Rubin, 2015; Li et al., 2018; Zhou et al., 2018; Li et al., 2020; Shi et al., 2022; Wang et al., 2023; Lu et al., 2023).
As the name of rerandomization indicates, researchers typically conduct rerandomization by repeatedly sampling treatment assignments until a balanced one is generated. This procedure is then repeated thousands of times since numerous independent qualified assignments are required to perform randomization-based inference. Therefore, rerandomization can be time-consuming using the acceptance-rejection sampling rerandomization(ARSRR) strategy. Zhu and Liu (2022) proposed a pair-switching rerandomization(PSRR) method to reduce the computational cost of rerandomization. However, the authors only considered a fixed number of covariates and a fixed covariate imbalance threshold. As we will show in Section 5, PSRR can be slower than ARSRR in certain settings. Moreover, both ARSRR and PSRR are computationally inefficient in settings considered by recent works, including diminishing covariate imbalance and high-dimensional covariates (Branson and Shao, 2021; Wang and Li, 2022; Zhang et al., 2023). Additionally, although Zhu and Liu (2022) extended PSRR to sequential experiments, PSRR is not applicable in other structured experiments such as stratified experiments or experiments with clusters.
In this work, we point out that a treatment assignment with a sufficiently small Mahalanobis distance can be achieved by optimizing a quadratic function of linearly constrained binary variables. Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to accelerate rerandomization in various randomized experiments. Moreover, we derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR. Our simulation studies and real data analysis illustrate that VNSRR significantly improves the computational efficiency of rerandomization while maintaining desired statistical properties. It can sample thousands of treatment assignments within seconds, even in cases where existing methods require hours to complete the task.
2 Related Work
2.1 The Neyman-Rubin Potential Outcome Framework
This study adopts the Neyman-Rubin potential outcome framework (Neyman, 1923; Rubin, 1974). We consider an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group. We denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators. $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise. Each unit $i$ has two potential outcomes, $(Y_{i}(1),Y_{i}(0))$, which correspond to the two treatment arms. The individual treatment effect for unit $i$ is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$. And the average treatment effect is defined as $\tau=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}(1)-Y_{i}(0))$. The observed outcome for unit $i$ is $Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0)$. We adopt the stable unit treatment value assumption(SUTVA) (Rubin, 1980), which states that the treatment levels are well-defined and there is no interference between $(Y_{i}(1),Y_{i}(0))$ and $\{W_{j}:j\neq i\}$. We estimate the treatment effects with the difference-in-means estimator, defined as $\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0)$. We assume that each unit has $p$ observed covariates, denoted by $X_{i}=(X_{i1},\ldots,X_{ip})^{T}$. The covariates of all units are gathered into a matrix $X=(X_{1},\ldots,X_{n})^{T}$. The corresponding covariance matrix is denoted by $S_{XX}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{T}$.
2.2 Rerandomization Using the Mahalanobis Distance
For a given assignment $W$, the difference in the covariate means of the two treatment groups is
$\overline{X}_{t}-\overline{X}_{c}$ $=\frac{1}{n_{t}}X^{T}W-\frac{1}{n_{c}}X^{T}(1_{n}-W)$
$=\frac{n}{n_{t}n_{c}}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
And the Mahalanobis distance between the covariates means is defined as
$M(W)$ $:=\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}\left[Cov[\left(\overline{X}_{t}-\overline{X}_{c}\right)]\right]^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
$=\frac{n_{t}n_{c}}{n}\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}S_{XX}^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
$=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n})$
$=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
where $H=\frac{n}{n_{t}n_{c}}XS_{XX}^{-1}X^{T}$.
Morgan and Rubin (2012) suggested to conduct rerandomization by sampling $W$ from $\{W:\sum_{i=1}^{n}W_{i}=n_{t},W_{i}\in\{0,1\}\}$ repeatedly until $M(W)\leq a$, where $a$ is a prespecified threshold. We refer to rerandomization based on this acceptance-rejection sampling strategy as ARSRR.
2.3 Pair-switching Rerandomization (PSRR)
To avoid the computationally inefficient acceptance-rejection sampling approach and accelerate rerandomization in both nonsequential and sequential experiments, Zhu and Liu (2022) suggested to search for balanced assignments with a pair-switching strategy, i.e., PSRR. The algorithm essentially starts with a random assignment and gradually improves it by switching between a treated unit and a control unit repeatedly. Motivated by the Metropolis-Hastings algorithm, PSRR allows moving to less balanced assignments with a non-zero probability to prevent being trapped in a local region. The authors obtained the unbiasedness and a lower bound on the variance reduction of the treatment effect estimator under PSRR. However, as pointed out in Section 1 and demonstrated by the simulation studies in Section 5, PSRR is still time-consuming in many settings and is not applicable in certain structured experiments.
3 Variable Neighborhood Searching Rerandomization(VNSRR)
We introduce our variable neighborhood searching rerandomization (VNSRR) in this section.
3.1 The VNSRR Algorithm
We notice that rerandomization is closely related with the following constrained binary quadratic programming problem.
minimize $\quad M(W)=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
subject to $\quad\sum_{i=1}^{n}W_{i}=n_{t},$
$\quad W_{i}\in\{0,1\},\;i=1,\ldots,n.$ (1)
For simplification, we denote the feasible set of (3.1) as $\mathcal{F}_{0}=\{W\in\{0,1\}^{n}:\sum_{i}W_{i}=n_{t}\}$. While rerandomization aims to find a feasible $W$ in $\mathcal{F}_{0}$ with sufficiently small $M(W)$ rather than the minimizer of $M(W)$ in $\mathcal{F}_{0}$, we can still utilize combinatorial optimization techniques to conduct rerandomization and are allowed to terminate the optimization process as long as $M(W)$ falls below a prespecified threshold $a$.
There are few studies concerning the specific optimization problem in (3.1), but some related constrained binary quadratic programming problems have received extensive attention, such as the maximum diversity problem and the heaviest k-subgraph problem. These problems share the same feasible sets as (3.1) but have different objective functions. We recommend that interested readers refer to Martí et al. (2013) for a detailed review of the literature. The authors presented extensive experiments to reveal that the best methods to solve the maximum diversity problem are local search based metaheuristics, especially variable neighborhood search (Brimberg et al., 2009; Aringhieri and Cordone, 2011) and tabu search (Palubeckis, 2007; Aringhieri and Cordone, 2011). Both methods update their solutions by exchanging coordinates of $W$ that are ones with coordinates of $W$ that are zeroes. Variable neighborhood search systematically changes the neighborhood by executing two steps iteratively. It descents to find local optimum in local search steps and perturbs to avoid being trapped in local regions in shaking steps. Tabu search maintains a list of forbidden moves, namely the tabu list, to avoid looping over recently visited coordinates of $W$. Aringhieri and Cordone (2011) proposed to incorporate tabu search in the local search step of variable neighborhood search.
Input: Covariates $X$, group sizes $(n_{t},n_{c})$, threshold $a$, hyperparameter $(L,S)$.
Set $t=0$;
Set $W^{(0)}$ as $n_{t}$ elements equal to 1 and $n_{c}$ elements equal to 0 with random positions;
Set $M^{(0)}=M(W^{(0)})$;
while $M^{(t)}>a$ do
       Randomly select $L$ positions of the 1’s in $W^{(t)}$, denoted by $O_{1},\cdots,O_{L}$;
       Randomly select $L$ positions of the 0’s in $W^{(t)}$, denoted by $Z_{1},\cdots,Z_{L}$;
       Set $\widetilde{S}=S$;
       for $l=1,\ldots,L$ do
             Obtain $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$;
             Set $M^{*}=M(W^{*})$;
             if $M^{*}<M^{(t)}$ then
                   Set $\widetilde{S}=0$;
                   Set $t=t+1$;
                   Set $W^{(t)}=W^{*}$;
                   Set $M^{(t)}=M^{*}$;
                   if $M^{(t)}<a$ then break;
      if $\widetilde{S}>0$ then
             Randomly select $\widetilde{S}$ positions of the 1’s in $W^{(t)}$, denoted by $O_{1},\cdots,O_{S}$;
             Randomly select $\widetilde{S}$ positions of the 0’s in $W^{(t)}$, denoted by $Z_{1},\cdots,Z_{S}$;
             for $s=1,\ldots,\widetilde{S}$ do
                   Obtain $W^{(t+1)}$ by switching the 1 at position $O_{s}$ and the 0 at position $Z_{s}$ in $W^{(t)}$;
                   Set $M^{(t+1)}=M(W^{(t+1)})$;
                   Set $t=t+1$;
Output: $W=W^{(t)}$.
Algorithm 1 Variable neighborhood searching rerandomization
Motivated by the empirical success of local-search-based metaheuristics in related optimization problems, we propose a novel VNSRR method for rerandomization through variable neighborhood searching, as summarized by Algorithm 1. VNSRR starts with a random assignment and then executes the local search step and the shaking step iteratively until an acceptable assignment is reached. VNSRR conducts the local search step differently from any existing variable neighborhood search methods. It is inspired by but different from tabu search methods. In the local search step, VNSRR sequentially examines $L$ disjoint random pairs of treated units and control units and swaps a pair if it leads to a more balanced assignment. We borrow the idea of preventing looping over recently visited units from tabu search. However, VNSRR differs from tabu search in that VNSRR avoids maintaining the tabu list and checking whether each swap is forbidden, further improving computational efficiency. We note that when selecting multiple treated units or control units in the local search step, the ordering of the units is also randomly selected. In other words, we randomly select a permutation of $L$ positions of the 1’s or 0’s in $W^{(t)}$. In the shaking step, VNSRR generates an assignment in the neighborhood of the current assignment by randomly switching $\widetilde{S}$ pairs of treated and control units. When no pairs are switched in the local search step, indicating a likely local optimum, $\widetilde{S}$ is set as a positive value $S$ to prevent getting stuck. Otherwise, $\widetilde{S}$ is set as 0, and no perturbation is performed to allow for further search in the current local region. We compare different choices of $L$ and $S$ in our experiments, and the results are relegated to the Supplement.
We point out that both PSRR and VNSRR use pairwise swap as the basic operation in their algorithms, but use it with different strategies. PSRR allows switching recently considered units, while VNSRR does not. PSRR, as motivated by simulated annealing and Metropolis-Hastings algorithm, accepts swaps that lead to worse covariate balance, i.e., larger $M(W)$, with a certain probability determined by a temperature parameter. In contrast, VNSRR keeps moving to new $W$ with smaller $M(W)$ in the local search step until none of the $L$ swaps lead to smaller $M(W)$, and then the shaking step is conducted. Our experiments illustrate that PSRR can move to $W$ with worse $M(W)$ quite frequently, while the strategy of VNSRR leads to a faster descent of $M(W)$. In fact, PSRR with zero temperature is equivalent to VNSRR with $L=1$ and without shaking steps. Our experiments show that this special version lies in the middle of PSRR and VNSRR. The zero-temperature version of PSRR was not considered by the authors of PSRR and is faster than their default version. Our default VNSRR is faster than VNSRR with $L=1$ and without shaking steps.
Like PSRR, we adopt the following trick to update the Mahalanobis distance efficiently. When we obtain a new assignment $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$, $M(W^{*})$ can be efficiently calculated as
$M(W^{*})=$ $M^{(t)}-2\sum_{i=1}^{n}W_{i}^{(t)}H_{O_{l}i}+H_{O_{l}O_{l}}+2\sum_{i=1}^{n}W_{i}^{*}H_{Z_{l}i}-H_{Z_{l}Z_{l}}+h_{O_{l}}-h_{Z_{l}},$
where $h=2\frac{n_{t}}{n}H1_{n}$.
We note that some deterministic or semi-deterministic algorithms might generate a single acceptable assignment faster than VNSRR. However, such algorithms cannot be used to generate numerous independent assignments and do not guarantee the unbiasedness of the corresponding treatment effect estimator. The next section discusses the statistical properties of different rerandomization methods.
3.2 Statistical Properties of VNSRR
In this section, we establish the statistical properties of various rerandomization algorithms based on the Mahalanobis distance. For clarification, a rerandomization algorithm here is a sampling algorithm that randomly generates a balanced treatment assignment from the feasible set $\mathcal{F}\subseteq\mathcal{F}_{0}$ by generating a chain of treatment assignments, $(W^{(0)},\ldots,W^{(T)})$, where $W^{(t)}\in\mathcal{F}$ for all $t$ and $W^{(T)}\in\mathcal{F}\cap\{W:M(W)\leq a\}$ is the final output. We provide sufficient conditions for a rerandomization algorithm to yield unbiased treatment effect estimators with bounded variance. Specifically, we consider the following conditions, which assumes Morgan and Rubin (2012) as well as Zhu and Liu (2022).
Condition 1
(linear decomposition) For $w_{i}=0,1$, $Y_{i}(w_{i})=\beta_{0}+\beta^{T}X_{i}+\tau w_{i}+e_{i}$, where $\beta_{0}+\beta^{T}X_{i}$ is the linear projection of $Y_{i}(0)$ onto $(1,X)$ and $e_{i}$ is the deviation from the linear projection.
Condition 2
(normality) $\widehat{\tau}$ and $\overline{X}_{t}-\overline{X}_{c}$ are normally distributed.
And we focus on three key properties of rerandomization algorithms.
Property 1
(symmetric feasible set) $\mathcal{F}=1-\mathcal{F}$.
Property 2
(uniform initialization) For any $w\in\mathcal{F}$,
$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
Property 3
(updating rule) There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
Based on Property 1-3, we prove that the difference-in-means estimator is unbiased.
Theorem 1
If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
Based on Condition 1-2 and Property 1-3, we prove that the difference-in-means estimator has smaller variance than that under complete randomization and construct a lower bound for the variance reduction. Denote $V_{CR}$ as the variance of the difference-in-means estimator under complete randomization, we prove the following theorem.
Theorem 2
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where $R^{2}=\frac{\beta^{T}Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]\beta}{V_{CR}}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$. Similar to the definition of $V_{CR}$, $Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]$ represents the covariance matrix of $\overline{X}_{t}-\overline{X}_{c}$ under complete randomization.
One can verify that VNSRR with $\mathcal{F}_{0}$ as the feasible set satisfies our proposed requirements under the following conditions.
Condition 3
(equal group sizes) $n_{t}=n_{c}=\frac{n}{2}$.
Proposition 1
VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
Therefore, the unbiasedness and variance reduction of the difference-in-means estimator under VNSRR are immediately evident as corollaries of Theorem 1 and Theorem 2.
Corollary 1
Under Condition 3 and VNSRR, $E[\widehat{\tau}]=\tau$.
Corollary 2
Under Condition 1, Condition 2, Condition 3, and VNSRR,
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
One can verify that CR, ARSRR, and PSRR all satisfy Property 1, Property 2, and Property 3 under certain conditions. As a result, their statistical properties are established by this framework as well. All verifications and proofs are included in the Supplement.
4 VNSRR for Structured Randomized Experiments
In this section, we extend VNSRR to randomized experiments with various types of structures. For each VNSRR variant, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator. We will reuse several notations across different settings, leveraging the similarities between settings to streamline our notation system. The pseudocode of each variant, along with the theoretical results and their proofs, is provided in the Supplement.
4.1 VNSRR for Sequentially Randomized Experiments
In Section 3, we consider offline rerandomization, which assigns treatment to all units at once. However, as pointed out by Zhou et al. (2018), experimenters may be unable to wait for all units to arrive before performing the experiments. It is common to assign treatment to units group by group in practice. To balance the covariates in such sequential experiments, Zhou et al. (2018) suggested performing sequential rerandomization, which can be computationally challenging. Bertsimas et al. (2019) proposed a novel algorithm for online allocation that leverages robust mixed-integer optimization. However, the authors focused on optimal designs, which are inherently different from the rerandomization framework.
We now review the basic concepts and notations of sequential rerandomization. We assume that the experiment consists of $K$ stages, and the $n$ units are divided into $K$ sequential groups of sizes $(n_{1},\ldots,n_{K})$. In the $k$-th stage, $n_{tk}$ units in the $k$-th group are assigned to the treatment arm, and the remaining $n_{ck}$ units are assigned to the control arm. We use the subscript $[k]$ to represent the reduced version of a term calculated based on only the first $k$ groups. The covariates of the first $k$ groups are denoted by $X_{[k]}$ and its covariance matrix is denoted by $S_{XX[k]}$. The assignment status of the first $k$ groups is denoted by $W_{[k]}$. And we define $n_{[k]}=\sum_{g=1}^{k}n_{g}$, $n_{t[k]}=\sum_{g=1}^{k}n_{tg}$, $n_{c[k]}=\sum_{g=1}^{k}n_{cg}$, $H_{[k]}=\frac{n_{[k]}}{n_{t[k]}n_{c[k]}}X_{[k]}S_{XX[k]}^{-1}X_{[k]}^{T}$. The Mahalanobis distance corresponding to $W_{[k]}$ is defined as
$M_{[k]}$ $=(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})^{T}H_{[k]}(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})$
Sequential rerandomization proceeds as follows: When the $k$-th group enrolls in the experiment, the experimenter randomly assigns treatment to the units without changing the assignment status of the previously enrolled groups. The experimenter aims to balance the covariates of all units in the first $k$ groups. Therefore, the experimenter rerandomizes the $n_{k}$ units until $M_{[k]}\leq a_{k}$, where $a_{k}$ is a prespecified threshold. Zhou et al. (2018) proposed to use acceptance-rejection sampling in this procedure, so we refer to their method as SeqARSRR. Recently, Zhu and Liu (2022) proposed a sequential pair-switching rerandomization(SeqPSRR) method to reduce the computational cost. SeqPSRR essentially performs PSRR in each stage of the sequential experiment. In Section 5.2, we will show that both SeqARSRR and SeqPSRR can be time-consuming.
We note that sequential rerandomization in the $k$-th stage can be formulated into the following constrained binary quadratic programming problem, whose feasible set we refer to as $\mathcal{F}_{[k]}$.
minimize $\quad M_{[k]}(W_{[k]}),$
subject to $\quad\sum_{i=n_{[k-1]}+1}^{n_{[k]}}W_{[k],i}=n_{tk},$
$\quad W_{[k],i}\in\{0,1\},\;i=n_{[k-1]}+1,\ldots,n_{[k]},$
$\quad W_{[k],i}=W_{[k-1],i},\;i=1,\ldots,n_{[k-1]}.$ (2)
Based on this observation, we extend VNSRR to sequentially randomized experiments and refer to this variant as SeqVNSRR. SeqVNSRR performs VNSRR in each stage of the experiment while fixing the assignments obtained in previous stages. It also adopts the updating trick, the specific form of the local search step, and the hyperparameter combination mentioned in Section 3 to reduce the computational cost.
To analyze the statistical properties of SeqVNSRR, we consider the following variants of Condition 3 and Property 1-3 in sequential experiments.
Condition 4
$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
Property 4
For any $k$, $\mathcal{F}_{[k]}=1-\mathcal{F}_{[k]}$.
Property 5
For any $k$ and $w\in\mathcal{F}_{[k]}$,
$P\left(W_{[k]}^{(0)}=w\right)=\frac{1}{|\mathcal{F}_{[k]}|}.$
Property 6
There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $k$, $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain for the $k$-th stage,
$P\left(W_{[k]}^{(t)}=w^{(t)}\mid W_{[k]}^{(0)}=w^{(0)},\ldots,W_{[k]}^{(t-1)}=w^{(t-1)}\right)=u_{t}(M_{[k]}(w^{(0)}),\ldots,M_{[k]}(w^{(t)})),$
Based on these conditions and properties, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under SeqVNSRR.
4.2 VNSRR for Stratified Randomized Experiments
Other than rerandomization, stratification stands out as another approach to balance covariates. The idea of stratification, or blocking, was first introduced by Fisher (1926) to balance discrete covariates. In stratified randomized experiments, experimenters divide units into strata, namely, layers or levels, according to the discrete covariates and conduct complete randomization within each stratum. Wang et al. (2023) proposed two strategies that combine stratification and rerandomization to improve covariate balance further. The overall strategy rerandomizes over all strata and measures the overall covariate imbalance, while the stratum-specific strategy rerandomizes within each strata separately and measures the stratum-specific covariate imbalance. Both strategies have their unique advantages. However, since the stratum-specific strategy is identical to the setting we consider in Section 3 within each stratum, we mainly focus on the overall strategy in this section.
We assume that the $n$ units are divided into $K$ strata of sizes $(n_{1},\ldots,n_{K})$. The index set of units in the $k$-th stratum is denoted by $\mathcal{I}_{k}$. In the $k$-th stratum, $n_{tk}$ units are assigned to the treatment arm and $n_{ck}=n_{k}-n_{tk}$ units are assigned to the control arm. Similar to Section 2, the assignment status is represented by a $n$-dimensional vector $W$ and the covariates are gathered into a $n\times p$ matrix $X$. The overall strategy proposed by Wang et al. (2023) samples assignments from the space $\{W\in\{0,1\}^{n}:\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\forall k\}$ repeatedly until $M(W)<a$. To our knowledge, the only existing method to achieve this goal is to perform acceptance-rejection sampling, which we refer to as StratARSRR.
We note that rerandomization in stratified experiments can be formulated into the following constrained binary quadratic programming problem.
minimize $\quad M(W),$
subject to $\quad\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\;k=1,\ldots,K$
$\quad W_{i}\in\{0,1\},\;i=1,\ldots,n.$ (3)
Based on this observation, we extend VNSRR to stratified randomized experiments and refer to this variant as StratVNSRR. StratVNSRR performs variable neighborhood searching while considering the constraints imposed by the stratum structure. In the local search step, StratVNSRR randomly selects a list of multiple disjoint pairs of treated units and control units from each stratum. Then StratVNSRR aggregates the lists from different strata and permutes the pairs in the aggregated list. StratVNSRR follows this new ordering of unit pairs to examine whether switching each pair leads to a more balanced assignment. Compared to consecutively examining pairs in the same stratum, this procedure avoids any predetermined ordering of units or strata and allows for more flexible search in the local region. In the shaking step, StratVNSRR conducts random unit swaps within every stratum. This ensures that the new assignment is sufficiently distant from the current one.
We note that the feasible set in (4.2) is a subset of the feasible set in (3.1). As a result, StratVNSRR belongs to the class of rerandomization methods we consider in Section 3. We verify that StratVNSRR always satisfies Property 2-3 and satisfies Property 1 under Condition 4. Therefore, the statistical properties of the difference-in-means estimator under StratVNSRR can be derived as corollaries of Theorem 1 and Theorem 2.
4.3 VNSRR for Cluster Randomized Experiments
Many public health and social science experiments (Donner et al., 2000; Hayes and Moulton, 2017; Schochet, 2020) assign the treatment at the cluster level rather than the individual level due to practical limitations or policy factors. Suppose the $n$ units are divided into $K$ clusters of sizes $(n_{1},\ldots,n_{K})$ in a cluster randomized experiment. We denote $\mathcal{I}_{k}$ as the index set of units in the $k$-th cluster. The experimenter assigns $K_{t}$ clusters to the treatment arm and $K_{c}=K-K_{t}$ clusters to the control arm. We note that when the $n_{i}$’s are not all equal, neither the treatment arm nor the control arm has a fixed number of units, which is different from all the settings we consider in previous sections. The assignment status of the clusters is denoted by a $K$-dimensional vector, $U$. And the assignment status of all units is denoted by a $n$-dimensional vector, $W$. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit.
Lu et al. (2023) suggested conducting cluster rerandomization to balance covariates in cluster randomized experiments. The authors proposed two metrics that measure the covariate imbalance. One is the Mahalanobis distance of cluster-level covariates, and the other is the Mahalanobis distance of individual-level covariates. Since using the first metric is equal to treating a cluster as an aggregated unit and can be efficiently solved by the approach proposed in Section 3, we mainly focus on the second metric in this section. We denote the individual-level covariates as a $n\times p$ matrix, $X$. Some cluster-level covariates may also be available, but we can construct individual-level covariates from cluster-level ones by ensuring they are identical within each cluster. We then formulate cluster rerandomization into the following constrained binary quadratic programming problem.
minimize $\quad M(W),$
subject to $\quad W_{i}=U_{k},\;\text{for all }k\text{ and }i\in\mathcal{I}_{k}$
$\quad\sum_{k=1}^{K}U_{k}=K_{t},$
$\quad U_{k}\in\{0,1\},\;k=1,\ldots,K.$ (4)
We propose the ClustVNSRR method, a variant of VNSRR, to solve (4.3) efficiently and accelerate cluster rerandomization. ClustVNSRR performs variable neighborhood searching at the cluster level and treats the clusters as the operational units in the local and shaking search steps. Similar to Section 4.2, we note that the feasible set in (4.3) is a subset of the feasible set in (3.1). Therefore, we derive the statistical properties of the difference-in-means estimator under ClustVNSRR by verifying that ClustVNSRR always satisfies Property 2-3 and satisfies Property 1 under the following variant of Condition 3.
Condition 5
$K_{t}=K_{c}=\frac{K}{2}$.
5 Simulation Studies and Real Applications
This section compares VNSRR with CR, ARSRR, and PSRR. To compare the computation time of VNSRR and the existing methods, we use each method to sample 1000 acceptable assignments respectively in one simulation and replicate the simulation 100 times on a MacBook Pro (Apple M1 Max Chip, 32 GB Memory, 10 Cores). To examine the statistical properties, we use each method to sample one acceptable assignment in the design stage to conduct the experiment and 1000 acceptable assignments in the analysis stage to conduct randomization-based inference. We replicate the simulation 1000 times using a High-Performance Computing cluster.
5.1 Simple Randomized Experiments
We first consider the simple randomized experiments described in Section 3 and mimic the simulation settings in Zhu and Liu (2022). We assume that the covariates follow the standard normal distribution identically and independently, $X_{ij}\overset{i.i.d.}{\sim}N(0,1)$, $i=1,\ldots,n$, $j=1,\ldots,p$. The potential outcomes of the control group are generated independently from a linear model, $Y_{i}(0)=\sum_{j=1}^{p}X_{ij}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,p)$. We conduct simulations under both the null hypothesis and the alternative hypothesis, where we set $Y_{i}(1)=Y_{i}(0)$ and $Y_{i}(1)=Y_{i}(0)+0.2\sqrt{Var[Y_{i}(0)]}$, respectively. We consider three sample sizes, $n\in\{30,100,500\}$. The size of the treatment group and the control group are set as equal.
Morgan and Rubin (2012) proved that the asymptotic acceptance probability of the assignment is $p_{a}=P(\chi_{p}^{2}<a)$ and Li et al. (2018) recommended that $p_{a}=10^{-3}$. Wang and Li (2022) considered settings where $p_{a}$ and $p$ are allowed to grow with $n$ rather than fixed. Motivated by these works, we considered two thresholds of the Mahalanobis distance, $p_{a}\in\{10^{-3},10^{-4}\}$, and two regimes for the number of covariates, $p=2$ and $p=n/2$.
We perform Fisher randomization tests with a significance level $\alpha=0.1$ and construct randomization-based confidence intervals with a nominal coverage rate of 90%. The exact formulas of Fisher randomization test p-values and randomization-based confidence intervals are provided in the Supplement. We use the default tuning parameter of PSRR, as Zhu and Liu (2022) pointed out that different tuning parameters lead to similar performance. For each method, we examine the absolute bias and standard deviation of the difference-in-means estimator, the size and power of the Fisher randomization test, the coverage and length of confidence intervals, and the average run time to sample $10^{3}$ acceptable assignments. We also use the randomness metric in Zhu and Liu (2022) to measure the randomness of different methods in the Supplement.
Table 1 shows the statistical and computational performances of all methods in two typical settings: a low-dimensional setting with $(n,p)=(30,2)$ and a high-dimensional setting with $(n,p)=(500,250)$. The complete results are provided in the Supplement. VNSRR is significantly faster than ARSRR and PSRR in all settings. VNSRR merely requires seconds to generate 1000 acceptable assignments, even in cases where ARSRR or PSRR requires an hour. For all methods, the finite-sample biases are negligible, Fisher randomization test controls the type-1 error, and the confidence intervals are valid. Compared to CR, all the rerandomization methods reduce the variance of the difference-in-means estimator, increase the statistical power, produce tighter confidence intervals, and have comparable performances in terms of these aspects. These results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and substantially reduces the computational cost in simple randomized experiments.
Table 1: Statistical and computational performance of different methods in simple randomized experiments.
$n$ $p$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
Bias SD Size Power Coverage Length
30 2 CR 6.2 74 10.5 14 89.5 246 $<0.1$
ARSRR 0.1 51 8.7 19 91.3 177 12.7
PSRR 1.5 51 10.1 18 89.9 177 50.3
VNSRR 1.6 52 11.0 17 89.0 177 1.4
500 250 CR 0.8 201 10.4 73 89.6 660 $<0.1$
ARSRR 4.0 186 9.2 77 90.8 619 4137.4
PSRR 4.5 189 9.8 78 90.2 619 275.3
VNSRR 7.6 199 12.6 73 87.4 615 1.2
5.2 Structured Randomized Experiments
This section considers the three types of structured experiments described in Section 4. We simulate the covariates and potential outcomes in the same way as in Section 5.1, except that we set $Y_{i}(1)=Y_{i}(0)+0.15\sqrt{Var[Y_{i}(0)]}$ under the alternative to ensure the power of different methods remains within a moderate range. For sequential experiments, we set the number of groups as $K=2$ and consider two group sizes, $n_{k}\in\{100,500\}$. We use the strategy proposed by Zhou et al. (2018) to determine $a_{k}$. Specifically, we set $(s_{1},s_{2})=(239,761)$ when $p=50$ and $(s_{1},s_{2})=(264,736)$ when $p=250$, where $a_{k}=\frac{n_{k}}{n_{[k]}}q_{k}$ and $q_{k}$ is the $\frac{1}{s_{k}}$ quantile of a non-central chi-square distribution with $p$ degrees of freedom and a non-central parameter $\frac{n_{[k-1]}}{n_{k}}M_{[k-1]}$. For stratified experiments, we set the stratum number as $K=2$ and consider two stratum sizes, $n_{k}\in\{100,500\}$. For cluster randomized experiments, we set the cluster size as $n_{k}=2$ and consider two options for the number of clusters, $K\in\{100,500\}$. We set $p_{a}=P(\chi_{p}^{2}<a)=10^{-3}$ for both stratified randomized experiments and cluster randomized experiments and set $p=n/4$ for all settings in this section.
Table 2 shows the results when $n=1000$. The complete results are provided in the Supplement. For clarity, we use a unified name to refer to different method variants. Since PSRR is not extended to stratified experiments and cluster experiments, it is excluded from the two schemes. The overall findings are consistent with the results in Section 5.1. VNSRR substantially reduces the computational burden of rerandomization, especially for high-dimensional datasets and in stratified randomized experiments and cluster randomized experiments, where ARSRR is the only method available to perform rerandomization. VNSRR maintains the statistical properties of rerandomization in all the settings and significantly improves statistical inference compared to CR.
Table 2: Statistical and computational performance of different methods in structured randomized experiments.
Scheme Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
Bias SD Size Power Coverage Length
Sequential CR 3.2 146 10.7 76 89.3 466 $<0.1$
ARSRR 6.0 139 12.2 80 87.8 428 3288.5
PSRR 3.3 135 11.5 82 88.5 429 339.1
VNSRR 5.8 138 11.4 81 88.6 428 5.8
Stratified CR 3.0 140 10.0 76 90.0 466 $<0.1$
ARSRR 2.7 131 9.1 82 90.9 434 3386.8
VNSRR 0.0 130 10.5 83 89.5 436 3.0
Cluster CR 0.9 140 9.8 77 90.2 465 $<0.1$
ARSRR 6.2 132 9.3 80 90.7 435 3116.2
VNSRR 3.3 131 10.1 82 89.9 436 3.9
5.3 A Clinical Trail Example
In addition to the simulation studies, we analyze a real dataset from a phase 1 clinical trial to illustrate the advantages of VNSRR. This dataset was analyzed by Zhu and Liu (2022) as well. The goal of this trial was to evaluate the interactions between intravenous methamphetamine and oral reserpine. In this trial, 20 participants were randomly assigned to the treatment group, and 10 participants were assigned to the control group. We aim to balance better three covariates, age, weight, and pre-treatment heart rate, which are significantly unbalanced in the original study, to better infer the treatment effects on post-treatment heart rate.
We set $a$ as the 0.001 quantile of $\chi_{3}^{2}$ and generate 10,000 acceptable assignments using each method. RR takes approximately 2.3 minutes to complete the task, and PSRR takes more than half an hour. In contrast, VNSRR requires only 12 seconds, which demonstrates that VNSRR accelerates the rerandomization process significantly. Based on the original dataset, we simulate semi-synthetic data and examine the statistical performance of each method. The results are provided in the Supplement.
6 Conclusion
In this work, we propose a novel VNSRR method to perform rerandomization efficiently in various types of randomized experiments. We derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under VNSRR. Our empirical results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and significantly reduces the computational cost compared to existing methods. A future direction is to extend the proposed method to experiments with tiers of covariates. Multi-objective optimization techniques are required to search for assignments that simultaneously balance the covariates within each tier.
Appendix
In Appendix A, we provide the pseudocode of the three VNSRR variants. In Appendix B, we state and prove all the theoretical results. In Appendix C, we introduce the Fisher randomization test and randomization-based confidence intervals. In Appendix D, we show the simulation results not included in the main text.
Appendix A Algorithmic Description of VNSRR Variants
A.1 Algorithmic Description of SeqVNSRR
The pseudocode of SeqVNSRR is presented in Algorithm 2. As mentioned in the main text, SeqVNSRR performs VNSRR in each experiment stage sequentially while fixing the assignments obtained in previous stages.
A.2 Algorithmic Description of StratVNSRR
The pseudocode of StratVNSRR is presented in Algorithm 3. StratVNSRR differs from VNSRR in that there are linear constraints within each stratum, $\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk}$. Therefore, StratVNSRR only exchanges units in the same stratum to ensure that the aforementioned linear constraints are always satisfied. StratVNSRR does not loop over the strata sequentially. Instead, after selecting the pairs of units to examine from each stratum, StratVNSRR aggregates the lists of pairs and randomly permutes the combined list. In this way, we ensure that all strata are equally treated, allowing for flexible exploration of the neighborhood without a predetermined ordering of the strata or the units.
A.3 Algorithmic Description of ClustVNSRR
The pseudocode of ClustVNSRR is presented in Algorithm 4. ClustVNSRR treats each cluster as a whole and uses the clusters as the operational units in the search process. In other words, ClustVNSRR essentially conducts VNSRR at the cluster level rather than the unit level. Recall that we denote the assignment status of the clusters as a $K$-dimensional vector $U$ and all units’ assignment status as a $n$-dimensional vector $W$. We have the following constraint imposed by the cluster structure of the experiment. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit. Since $W$ is determined once we generate $U$, we denote $W(U)$ as the mapping from $U$ to $W$ in Algorithm 4.
Input: Covariates $X$, number of units to assign to the two arms in each stage $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, thresholds $(a_{1},\ldots,a_{K})$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$.
for $k=1,\ldots,K$ do
       Set $t=0$;
       Initialize $W_{[k]}^{(0)}$: keep the elements obtained in previous stages fixed in $W_{[k]}^{(0)}$ if $k\geq 2$ and set the undetermined elements in $W_{[k]}^{(0)}$ as $n_{tk}$ elements equal to 1 and $n_{ck}$ elements equal to 0 with random positions;
       Set $M_{[k]}^{(0)}=M_{[k]}(W_{[k]}^{(0)})$;
       while $M_{[k]}^{(t)}>a_{k}$ do
             Randomly select $L_{k}$ positions of the 1’s in $W_{[k]}^{(t)}$, denoted by $O_{1},\cdots,O_{L_{k}}$;
             Randomly select $L_{k}$ positions of the 0’s in $W_{[k]}^{(t)}$, denoted by $Z_{1},\cdots,Z_{L_{k}}$;
             Set $\widetilde{S}=S_{k}$;
             for $l=1,\ldots,L_{k}$ do
                   Obtain $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W_{[k]}^{(t)}$;
                   Set $M^{*}=M(W^{*})$;
                   if $M^{*}<M^{(t)}$ then
                         Set $\widetilde{S}=0$;
                         Set $t=t+1$;
                         Set $W_{[k]}^{(t)}=W^{*}$;
                         Set $M_{[k]}^{(t)}=M^{*}$;
                         if $M_{[k]}^{(t)}<a_{k}$ then break;
            if $\widetilde{S}>0$ then
                   Randomly select $\widetilde{S}$ positions of the 1’s in $W_{[k]}^{(t)}$, denoted by $O_{1},\cdots,O_{S_{k}}$;
                   Randomly select $\widetilde{S}$ positions of the 0’s in $W_{[k]}^{(t)}$, denoted by $Z_{1},\cdots,Z_{S_{k}}$;
                   for $s=1,\ldots,\widetilde{S}$ do
                         Obtain $W_{[k]}^{(t+1)}$ by the 1 at positon $O_{s}$ and the 0 at positon $Z_{s}$ in $W_{[k]}^{(t)}$;
                         Set $M_{[k]}^{(t+1)}=M_{[k]}(W_{[k]}^{(t+1)})$;
                         Set $t=t+1$;
Output: $W=W^{(t)}$.
Algorithm 2 Sequential Variable neighborhood searching rerandomization
Input: Covariates $X$, index sets of units in each stratum $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of units to assign to the two arms in each stratum $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, threshold $a$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$, $L_{total}=\sum_{k}L_{k}$, $S_{total}=\sum_{k}S_{k}$.
Set $t=0$;
Set $W^{(0)}$ as $n_{tk}$ elements in $\mathcal{I}_{k}$ equal to 1 and $n_{ck}$ elements in $\mathcal{I}_{k}$ equal to 0 with random positions for all $k\in\{1,\ldots,K\}$;
Set $M^{(0)}=M(W^{(0)})$;
while $M^{(t)}>a$ do
       for $k=1,\ldots,K$ do
             Randomly select $L_{k}$ positions of the 1’s in $W^{(t)}$ that belong to $\mathcal{I}_{k}$, denoted by $O_{k1},\cdots,O_{kL_{k}}$;
             Randomly select $L_{k}$ positions of the 0’s in $W^{(t)}$ that belong to $\mathcal{I}_{k}$, denoted by $Z_{k1},\cdots,Z_{kL_{k}}$;
      Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{L_{total}})$ with random ordering;
       Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{L_{total}})$. The ordering of $(Z_{1},\ldots,Z_{L_{total}})$ is consistent with that of $(O_{1},\ldots,O_{L_{total}})$;
       Set $\widetilde{S}=S_{total}$;
       for $l=1,\ldots,L_{total}$ do
             Obtain $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$;
             Set $M^{*}=M(W^{*})$;
             if $M^{*}<M^{(t)}$ then
                   Set $\widetilde{S}=0$;
                   Set $t=t+1$;
                   Set $W^{(t)}=W^{*}$;
                   Set $M^{(t)}=M^{*}$;
                   if $M^{(t)}<a$ then break;
      if $\widetilde{S}>0$ then
             for $k=1,\ldots,K$ do
                   Randomly select $S_{k}$ positions of the 1’s in $W^{(t)}$ that are in $\mathcal{I}_{k}$, denoted by $O_{k1},\cdots,O_{kS_{k}}$;
                   Randomly select $S_{k}$ positions of the 0’s in $W^{(t)}$ that are in $\mathcal{I}_{k}$, denoted by $Z_{k1},\cdots,Z_{kS_{k}}$;
            Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{S_{total}})$;
             Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{S_{total}})$;
             for $s=1,\ldots,S_{total}$ do
                   Obtain $W^{(t+1)}$ by switching the 1 at position $O_{s}$ and the 0 at position $Z_{s}$ in $W^{(t)}$;
                   Set $M^{(t+1)}=M(W^{(t+1)})$;
                   Set $t=t+1$;
Output: $W=W^{(t)}$.
Algorithm 3 Stratified Variable neighborhood searching rerandomization
Input: Covariates $X$, index sets of units in each cluster $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of clusters in the treatment arm and the control arm $(K_{t},K_{c})$, threshold $a$, hyperparameter $(L,S)$.
Set $t=0$;
Set $U^{(0)}$ as $K_{t}$ elements equal to 1 and $K_{c}$ elements equal to 0 with random positions;
Set $M^{(0)}=M(W(U^{(0)}))$;
while $M^{(t)}>a$ do
       Randomly select $L$ positions of the 1’s in $U^{(t)}$, denoted by $O_{1},\cdots,O_{L}$;
       Randomly select $L$ positions of the 0’s in $U^{(t)}$, denoted by $Z_{1},\cdots,Z_{L}$;
       Set $\widetilde{S}=S$;
       for $l=1,\ldots,L$ do
             Obtain $U^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $U^{(t)}$;
             Set $M^{*}=M(W(U^{*}))$;
             if $M^{*}<M^{(t)}$ then
                   Set $\widetilde{S}=0$;
                   Set $t=t+1$;
                   Set $W^{(t)}=W^{*}$;
                   Set $U^{(t)}=U^{*}$;
                   if $M^{(t)}<a$ then break;
      if $\widetilde{S}>0$ then
             Randomly select $\widetilde{S}$ positions of the 1’s in $U^{(t)}$, denoted by $O_{1},\cdots,O_{S}$;
             Randomly select $\widetilde{S}$ positions of the 0’s in $U^{(t)}$, denoted by $Z_{1},\cdots,Z_{S}$;
             for $s=1,\ldots,\widetilde{S}$ do
                   Obtain $U^{(t+1)}$ by switching the 1 at position $O_{s}$ and the 0 at position $Z_{s}$ in $U^{(t)}$;
                   Set $M^{(t+1)}=M(W(U^{(t+1)}))$;
                   Set $t=t+1$;
Output: $W=W(U^{(t)})$.
Algorithm 4 Cluster Variable neighborhood searching rerandomization
Appendix B Additional Theoretical Results and Proofs
VNSRR, StratVNSRR, and ClustVNSRR all assign treatment in an offline manner. In contrast, SeqVNSRR performs online allocation and needs to be handled differently. We will first establish the theoretical properties of VNSRR, StratVNSRR, and ClustVNSRR, and then obtain the results of SeqVNSRR.
For the convenience of readers, we first restate the conditions we propose in the main text, which are related to the offline methods.
Condition 1
For $w_{i}=0,1$, $Y_{i}(w_{i})=\beta_{0}+\beta^{T}X_{i}+\tau w_{i}+e_{i}$, where $\beta_{0}+\beta^{T}X_{i}$ is the linear projection of $Y_{i}(0)$ onto $(1,X)$ and $e_{i}$ is the deviation from the linear projection.
Condition 2
$\widehat{\tau}$ and $\overline{X}_{t}-\overline{X}_{c}$ are normally distributed.
Condition 3
$n_{t}=n_{c}=\frac{n}{2}$.
Condition 4
$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
Condition 5
$K_{t}=K_{c}=\frac{K}{2}$.
Property 1
$\mathcal{F}=1-\mathcal{F}$.
Property 2
For any $w\in\mathcal{F}$,
$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
Property 3
There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
In this section, we first prove Theorem 1 and Theorem 2 based on the abovementioned conditions and properties.
Theorem 1
If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
Theorem 2
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where $R^{2}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$.
We then verify that VNSRR, StratVNSRR, and ClustVNSRR satisfy the requirements by proving the following propositions. Their unbiasedness and variance reduction are immediately evident as corollaries of Theorem 1 and Theorem 2.
Proposition 1
VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
Proposition 2
StratVNSRR satisfies Property 2 and Property 3. Under Condition 4, StratVNSRR satisfies Property 1.
Proposition 3
ClustVNSRR satisfies Property 2 and Property 3. Under Condition 5, ClustVNSRR satisfies Property 1.
We also show that the statistical properties of CR, ARSRR, and PSRR can be established in this way as well based on the following propositions.
Proposition 4
CR satisfies Property 1, Property 2 and Property 3.
Proposition 5
ARSRR satisfies Property 2 and Property 3. Under Condition 3, ARSRR satisfies Property 1.
Proposition 6
PSRR satisfies Property 2 and Property 3. Under Condition 3, PSRR satisfies Property 1.
Consequently, we will establish the statistical properties of SeqVNSRR by extending the conditions and properties we consider in the offline setting to the online setting.
B.1 Proof of Theorem 1
For any chain $\left(w^{(0)},\ldots,w^{(T)}\right)$ that the rerandomization algorithm may generate, we have
$P\left(W^{(0)}=w^{(0)}\right)=P\left(W^{(0)}=1-w^{(0)}\right),$
since the initial point $W(0)$ is sampled uniformly from a symmetric space $\mathcal{F}$.
If we exchange the assignments to the two arms by replacing $W$ with $1-W$, the covariate imbalance between the two arms should remain unchanged. In other words, $M(W)=M(1-W)$. As a result,
$\hskip 15.0ptP\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)$
$=u_{t}\left(M(w^{(0)}),\ldots,M(w^{(t)})\right)$
$=u_{t}\left(M(1-w^{(0)}),\ldots,M(1-w^{(t)})\right)$
$=P\left(W^{(t)}=1-w^{(t)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(t-1)}=1-w^{(t-1)}\right),$
Therefore, for any $w$, we have
$\hskip 15.0ptP\left(W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)},W^{(T)}=w^{(T)}\right)$
$=P\left(W^{(0)}=w^{(0)}\right)P\left(W^{(1)}=w^{(1)}\mid W^{(0)}=w^{(0)}\right)\cdots P\left(W^{(T)}=w^{(T)}\mid W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)}\right)$
$=P\left(W^{(0)}=1-w^{(0)}\right)P\left(W^{(1)}=1-w^{(1)}\mid W^{(0)}=1-w^{(0)}\right)\cdots$
$\hskip 15.0ptP\left(W^{(T)}=1-w^{(T)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)}\right)$
$=P\left(W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)},W^{(T)}=1-w^{(T)}\right),$
We then sum over all possible $\left(w^{(0)},\ldots,w^{(T-1)}\right)$ and obtain $P(W=w^{(T)})=P(W=1-w^{(T)})$. As a result, $P\left(W_{i}=1\right)=P\left(W_{i}=0\right)=1/2$. And we have
$E[\widehat{\tau}(W)]$ $=E\left[\frac{2}{n}\sum_{i=1}^{n}W_{i}Y_{i}(1)-\frac{2}{n}\sum_{i=1}^{n}\left(1-W_{i}\right)Y_{i}(0)\right]$
$=\frac{2}{n}\sum_{i=1}^{n}Y_{i}(1)P\left(W_{i}=1\right)-\frac{2}{n}\sum_{i=1}^{n}Y_{i}(0)P\left(W_{i}=0\right)$
$=\tau.$
B.2 Proof of Theorem 2
The difference-in-means estimator can be expressed as
$\widehat{\tau}$ $=\tau+\beta^{T}\left(\overline{X}_{t}-\overline{X}_{c}\right)+\left(\overline{e}_{t}-\overline{e}_{c}\right).$
$\overline{X}_{t}-\overline{X}_{c}$ and $\overline{e}_{t}-\overline{e}_{c}$ are uncorrelated since $\beta_{0}+\beta^{\mathrm{T}}X_{i}$ is the projection of $Y_{i}(0)$ onto $(1,X)$. In addition, $\widehat{\tau}$ and $\overline{X}_{t}-\overline{X}_{c}$ are normally distributed, so $\overline{X}_{t}-\overline{X}_{c}$ and $\overline{e}_{t}-\overline{e}_{c}$ are independent. Since the rerandomization algorithm only affects $\overline{X}_{t}-\overline{X}_{c}$ and does not affect $\overline{e}_{t}-\overline{e}_{c}$, we have
$Var[\widehat{\tau}]$ $=Cov\left[\beta^{T}(\overline{X}_{t}-\overline{X}_{c})\beta\right]+Var\left[\overline{e}_{t}-\overline{e}_{c}\right]$
$=\beta^{T}Cov\left[\overline{X}_{t}-\overline{X}_{c}\right]\beta+\left(1-R^{2}\right)V_{CR}$
We then bound the first term in the variance of $\widehat{\tau}$.
Recall that
$M=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
We define
$Z=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}(W-\frac{n_{t}}{n}1_{n})=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}(\overline{X}_{t}-\overline{X}_{c}).$
Then we have $M=Z^{T}Z$.
$E[Z]=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}E[W-\frac{1}{2}1_{n}]=0$. Therefore, $E[Z_{j}]=0$.
We note that the rows of $S_{XX}^{-1/2}X^{T}$ are exchangeable since the covariates are normalized. As a result, $\{Z_{j}\}_{1\leq j\leq p}$ are exchangeable, and we have
$Var[Z_{j}]=E[Z_{j}^{2}]=\frac{1}{p}E\left[\sum_{j=1}^{p}Z_{j}^{2}\right]=\frac{1}{p}E[M]\leq\frac{a}{p}.$
If we flip the sign of the $j$-th normalized covariate, the sign of $Z_{j}$ is also flipped, but the mapping $M(W)$ remains unchanged. Since the updating rule only depends on the Mahalanobis distance, this operation does not affect the distribution of $W$, hence the distribution of $Z_{j}$. As a result, $(Z_{j},Z_{l})$ and $(-Z_{j},Z_{l})$ are identically distributed. And we have
$Cov[Z_{j},Z_{l}]$ $=E[Z_{j}Z_{l}]$
$=E[E\left[Z_{j}Z_{l}\mid Z_{l}\right]]$
$=E[Z_{l}E\left[Z_{j}\mid Z_{l}\right]]$
$=E[Z_{j}\times 0]$
$=0.$
Therefore, $Cov[Z]=Var[Z_{j}]I_{p}$, and
$\beta^{T}Cov[\overline{X}_{t}-\overline{X}_{c}]\beta$ $=\beta^{T}Cov\left[\sqrt{\frac{n_{t}n_{c}}{n}}S_{XX}^{1/2}Z\right]\beta$
$=\frac{n_{t}n_{c}}{n}\beta^{T}S_{XX}^{1/2}Cov[Z]S_{XX}^{1/2}\beta$
$=\frac{n_{t}n_{c}}{n}Var[Z_{j}]\beta^{T}S_{XX}\beta$
$=Var[Z_{j}]R^{2}V_{CR}$
$\leq\frac{a}{p}R^{2}V_{CR}.$
Consequently,
$Var[\widehat{\tau}]\leq\frac{a}{p}R^{2}V_{CR}+(1-R^{2})V_{CR}.$
Rearranging the terms, we obtain
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
B.3 Proof of the Propositions
Recall that $\mathcal{F}_{0}=\{W\in\{0,1\}^{n}:\sum_{i}W_{i}=n_{t}\}$. Under Condition 3, for any $W\in\mathcal{F}_{0}$,
$\sum_{i}(1-W_{i})=n-n_{t}=n_{c}=n_{t}.$
This implies that $1-W\in\mathcal{F}_{0}$ as well. Thus, $\mathcal{F}_{0}=1-\mathcal{F}_{0}$. Since CR, ARSRR, PSRR, and VNSRR treat $\mathcal{F}_{0}$ as the feasible set, they all satisfy Property 1.
StratVNSRR and ClustVNSRR conduct rerandomization with more constraints imposed by the structure of the experiments. We show that these constraints are also symmetric in nature.
We denote the feasible set of StratVNSRR and ClustVNSRR as $\mathcal{F}_{Strat}$ and $\mathcal{F}_{Clust}$, respectively.
Under Condition 4, for any $W\in\mathcal{F}_{Strat}$,
$\sum_{i\in\mathcal{I}_{k}}(1-W_{i})=n_{k}-n_{tk}=n_{ck}=n_{tk}.$
This implies that $1-W\in\mathcal{F}_{Strat}$ and $\mathcal{F}_{Strat}=1-\mathcal{F}_{Strat}$.
Under Condition 5, for any $W\in\mathcal{F}_{Clust}$, suppose $U$ represents the assignment status of the clusters that corresponds to $W$. Then $1-U$ corresponds to $1-W$. And we have
$\sum_{k=1}^{K}(1-U_{k})=K-K_{t}=K_{c}=K_{t}.$
This implies that $1-W\in\mathcal{F}_{Clust}$ and $\mathcal{F}_{Clust}=1-\mathcal{F}_{Clust}$.
Therefore, both StratVNSRR and ClustVNSRR satisfy Property 1.
All of the methods satisfy Property 2 since they all start with a random assignment by uniformly sampling from the feasible set.
Regarding the updating rule, CR automatically satisfies Property 3 since it does not update the treatment assignment.
For ARSRR, we have
$u_{t}=\frac{1}{|\mathcal{F}|.}$
For PSRR, we have
$u_{t}=\left(\frac{M(w^{(t-1)})}{M(w^{(t)})}\right)^{\gamma}.$
Therefore, both ARSRR and PSRR satisfy Property 3.
For VNSRR and its variants, $(M(w^{(0)}),\ldots,M(w^{(t-1)}))$ determines whether $W^{(t)}$ is obtained through the local search step or the shaking step.
In the local search step, we have
$u_{t}=1\{M(w^{(t)})<M(w^{(t-1)})\}.$
And in the shaking step, $u_{t}$ is a constant which depends on the number of adjacent feasible solutions.
Therefore, VNSRR satisfies Property 3.
B.4 Theoretical Results for SeqVNSRR
In sequential experiments, we have the following Condition 4, which corresponds to the Condition 3 in the non-sequential settings.
Condition 4
$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
We also have the corresponding version of Property 1-3, stated as follows.
Property 4
For any $k$, $\mathcal{F}_{[k]}=1-\mathcal{F}_{[k]}$.
Property 5
For any $k$ and $w\in\mathcal{F}_{[k]}$,
$P\left(W_{[k]}^{(0)}=w\right)=\frac{1}{|\mathcal{F}_{[k]}|}.$
Property 6
There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $k$, $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain for the $k$-th stage,
$P\left(W_{[k]}^{(t)}=w^{(t)}\mid W_{[k]}^{(0)}=w^{(0)},\ldots,W_{[k]}^{(t-1)}=w^{(t-1)}\right)=u_{t}(M_{[k]}(w^{(0)}),\ldots,M_{[k]}(w^{(t)})),$
Based on these sufficient conditions, we obtain that
Theorem 3
If $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then $E[\widehat{\tau}]=\tau$.
Theorem 4
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a_{K}}{p}\right)R^{2},$
We omit the proofs of Theorem 3 and Theorem 4, since they are almost the same as the proofs of Theorem 1 and Theorem 2, except that the chain $(W^{(0)},\ldots,W^{(T)})$ needs to be decomposed into $K$ parts corresponding to the $K$ stages. Each part can be analyzed in the same way as in simple randomized experiments.
We have verified that VNSRR satisfies Property 1-3. We note that SeqVNSRR can be viewed as a special version of VNSRR in each of its stages. As a result, SeqVNSRR satisfies the sequential version of Property 1-3.
Proposition 7
SeqVNSRR satisfies Property 5 and Property 6. Under Condition 4, SeqVNSRR satisfies Property 4.
The unbiasedness and variance reduction properties of SeqVNSRR are then immediately evident as corollaries of Theorem 3 and Theorem 4.
Appendix C Randomization-based Inference
We use the Fisher randomization test to test the sharp null hypothesis, $H_{0}:Y_{i}(1)-Y_{i}(0)=0,i=1,\ldots,n$, versus the two-sided alternative hypothesis, $H_{1}:Y_{i}(1)-Y_{i}(0)\neq 0,i=1,\ldots,n$. As mentioned in Section 2, the observed outcome for unit $i$ is $Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0)$. Therefore, under $H_{0}$, $Y_{i}(1)=Y_{i}(0)=Y_{i}$, and we can impute all the potential outcomes with the observed outcomes. We then know the distribution of $\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0)$ since we can calculate $\widehat{\tau}(W)$ for every acceptable $W$. However, it is unrealistic to calculate the exact distribution of $\widehat{\tau}(W)$ due to the large size of the feasible set. As a result, we approximate the distribution of $\widehat{\tau}(W)$ using the Monte Carlo method. We sample $B=1000$ acceptable assignments $\{W^{b}\}_{1\leq b\leq B}$ independently and calculate $\{\widehat{\tau}(W^{b})\}_{1\leq b\leq B}$. The p-value is then calculated by comparing $\{\widehat{\tau}(W^{b})\}_{1\leq b\leq B}$ with the observed treatment effect estimator $\widehat{\tau}(W^{obs})$, $pv=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b})|\geq|\widehat{\tau}(W^{obs})|\}$
As pointed out by Zhu and Liu (2022), randomization-based confidence intervals can be constructed by inverting Fisher randomization tests. To construct a lower confidence bound for $\tau$, we consider testing $H_{0}:Y_{i}(1)-Y_{i}(0)=\theta,i=1,\ldots,n$, versus $H_{1}:Y_{i}(1)-Y_{i}(0)>\theta,i=1,\ldots,n$. Similar to testing the sharp null hypothesis with $\theta=0$, we can impute all the potential outcomes under $H_{0}$. We sample $B=1000$ acceptable assignments $\{W^{b}\}_{1\leq b\leq B}$ independently and calculate the corresponding treatment effect estimators $\{\widehat{\tau}(W^{b},\theta)\}_{1\leq b\leq B}$ using the imputed potential outcomes. The p-value is then calculated as $pv(\theta)=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b},\theta)|\geq|\widehat{\tau}(W^{obs})|\}$. Luo et al. (2021) proved that $\theta_{l}=\sup\{\theta:pv(\theta)\leq\alpha\}$ is a lower confidence bound for $\tau$ with confidence level $\alpha$. And Zhu and Liu (2022) proposed an efficient approach to solve for $\theta_{l}$, which we adopt in our numerical experiments. Similarly, we can construct an upper confidence bound $\theta_{u}$ with confidence level $\alpha$. $[\theta_{l},\theta_{u}]$ is then a confidence interval for $\tau$ with confidence level $1-2\alpha$.
Appendix D Simulation Results
In Section D.1, we provide the complete results of the simulation studies in the main text. In Section D.2, we examine the randomness of assignments generated by each method. In Section D.3, we generate semi-synthetic data based on the clinical trial dataset and compare the statistical performance of each method.
D.1 Statistical and Computational Performances
Table 3 and Table 4 show the complete results of our simulation studies for simple experiments and structured experiments, respectively. The results of ARSRR in the last setting of Table 3 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. The overall findings are similar to the conclusion in the main text. VNSRR substantially reduces the computational burden of rerandomization and maintains the appealing statistical properties of rerandomization.
D.2 Randomness Metric
Similar to Zhu and Liu (2022), we denote $L_{n}$ as the largest eigenvalue of the covariance matrix of $12W-$ 1 and use $L_{n}$ to measure the randomness of assignments generated by each method. As pointed out by Zhu and Liu (2022), larger $L_{n}$ implies less randomness.
Table 5 and Table 6 show the results in simple experiments and structured experiments, respectively. Similar to Table 3, the results of ARSRR in the last setting of Table 5 where $(n,p,p_{a})=(500,250,10^{-4})$ are omitted due to the substantial computational burden. CR is the most random method as expected. The three rerandomization methods perform similarly in all settings. Therefore, one can safely perform randomization-based statistical inference based on the assignments generated by VNSRR without the risk of relying on a small set of unique assignments.
D.3 Semi-synthetic Data
We mimic the set-ups in Zhu and Liu (2022) and simulate semi-synthetic data based on the clinical trial dataset. Since we only observe half of the potential outcomes in the original dataset, we impute the counterfactual potential outcomes under the sharp null. We then compare VNSRR and the existing methods in the same way as in our simulation studies. The results are shown in Table 7 and are similar to our findings in the simulation studies. VNSRR significantly accelerates rerandomization and has comparable performances in other aspects compared to ARSRR and PSRR.
Table 3: Statistical and computational performance of different methods in simple randomized experiments.
$n$ $p$ $p_{a}$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
Bias SD Size Power Coverage Length
30 2 $10^{-3}$ CR 6.2 74 10.5 14 89.5 246 $<0.1$
ARSRR 0.1 51 8.7 19 91.3 177 11.4
PSRR 1.5 51 10.1 18 89.9 177 48.1
VNSRR 1.6 52 11.0 17 89.0 177 1.4
15 $10^{-3}$ CR 9.2 199 9.3 18 90.7 678 $<0.1$
ARSRR 2.2 155 10.8 21 89.2 552 64.2
PSRR 3.8 156 10.6 22 89.4 547 1.6
VNSRR 6.8 152 10.4 22 89.6 547 0.3
100 2 $10^{-3}$ CR 1.4 42 11.7 29 88.3 132 $<0.1$
ARSRR 0.2 29 10.4 40 89.6 94 14.0
PSRR 0.3 29 9.6 40 90.4 94 2.1
VNSRR 0.4 27 8.5 40 91.5 94 0.5
$10^{-4}$ CR 1.4 42 11.7 29 88.3 132 $<0.1$
ARSRR 0.0 28 9.8 40 90.2 94 139.6
PSRR 0.1 28 10.1 40 89.9 94 35.7
VNSRR 0.3 28 10.2 41 89.8 94 4.2
50 $10^{-3}$ CR 1.0 207 11.3 35 88.7 663 $<0.1$
ARSRR 1.4 174 9.9 39 90.1 579 215.3
PSRR 5.1 170 10.1 42 89.9 574 8.2
VNSRR 2.1 173 10.1 40 89.9 573 0.2
$10^{-4}$ CR 1.0 207 11.3 35 88.7 663 $<0.1$
ARSRR 10.2 165 7.8 39 92.2 567 2453.3
PSRR 2.8 167 9.9 43 90.1 560 27.0
VNSRR 6.4 163 9.1 42 90.9 558 0.3
500 2 $10^{-3}$ CR 1.0 18 8.7 74 91.3 59 $<0.1$
ARSRR 0.4 13 8.8 94 91.2 42 24.3
PSRR 0.6 12 9.1 94 90.9 42 1.6
VNSRR 0.1 13 10.5 93 89.5 42 0.5
$10^{-4}$ CR 1.0 18 8.7 74 91.3 59 $<0.1$
ARSRR 0.5 12 0.2 95 90.8 42 248.8
PSRR 0.5 13 11.2 92 88.8 42 6.4
VNSRR 0.2 12 11.3 92 88.7 42 1.7
250 $10^{-3}$ CR 0.8 201 10.4 73 89.6 660 $<0.1$
ARSRR 4.0 186 9.2 77 90.8 619 4137.4
PSRR 4.5 189 9.8 78 90.2 619 275.3
VNSRR 7.6 199 12.6 73 87.4 615 1.2
$10^{-4}$ CR 0.8 201 10.4 88 89.6 660 $<0.1$
PSRR 4.3 187 10.3 91 89.7 611 2878.8
VNSRR 5.7 181 9.8 92 90.2 608 1.3
Table 4: Statistical and computational performance of different methods in structured randomized experiments.
Scheme $n$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
Bias SD Size Power Coverage Length
Sequential 200 CR 5.5 140 10.3 28 89.7 469 $<0.1$
ARSRR 3.3 112 8.1 36 91.9 391 162.4
PSRR 1.7 118 10.4 36 89.6 391 14.1
VNSRR 0.9 116 10.2 35 89.8 390 0.9
1000 CR 3.2 146 10.7 76 89.3 466 $<0.1$
ARSRR 6.0 139 12.2 80 87.8 428 3288.5
PSRR 3.3 135 11.5 82 88.5 429 339.1
VNSRR 5.8 138 11.4 81 88.6 428 5.8
Stratified 200 CR 5.2 140 10.3 26 89.7 467 $<0.1$
ARSRR 4.5 121 10.6 35 89.4 401 148.0
VNSRR 6.1 124 10.3 32 89.7 402 0.5
1000 CR 3.0 140 10.0 76 90.0 466 $<0.1$
ARSRR 2.7 131 9.1 82 90.9 434 3386.8
VNSRR 0.0 130 10.5 83 89.5 436 3.0
Cluster 200 CR 3.0 141 10.9 27 89.1 479 $<0.1$
ARSRR 3.3 116 8.8 32 91.2 405 126.4
VNSRR 6.4 116 8.9 30 91.1 405 0.8
1000 CR 0.9 140 9.8 77 90.2 465 $<0.1$
ARSRR 6.2 132 9.3 80 90.7 435 3116.2
VNSRR 3.3 131 10.1 82 89.9 436 3.9
Table 5: $L_{n}$ (×$\times 10^{-2}$) of different methods in simple randomized experiments.
$n$ $p$ $p_{a}$ Method
CR ARSRR PSRR VNSRR
30 2 $10^{-3}$ 137 158 161 158
15 $10^{-3}$ 137 347 333 329
100 2 $10^{-3}$ 170 174 174 174
$10^{-4}$ 170 174 174 174
50 $10^{-3}$ 170 231 234 233
$10^{-4}$ 170 242 248 245
500 2 $10^{-3}$ 288 289 289 289
$10^{-4}$ 288 289 289 289
250 $10^{-3}$ 288 307 307 310
$10^{-4}$ 288 NA 314 317
Table 6: $L_{n}$ (×$\times 10^{-2}$) of different methods in structured randomized experiments.
Scheme $n$ Method
CR ARSRR PSRR VNSRR
Sequential 200 208 251 252 253
1000 396 411 410 412
Stratified 200 208 228 NA 228
1000 396 404 NA 404
Cluster 200 339 438 NA 430
1000 576 606 NA 604
Table 7: Statistical and computational performance of different methods when applied to semi-synthetic data.
Method Statistical Inference (×$\times 10^{-3}$) Computation Time (second) $L_{n}$
Bias SD Size Coverage Length
CR 9.7 440 4.4 95.6 1804 $<0.1$ 122
ARSRR 2.9 317 4.3 95.7 1333 13.9 140
PSRR 0.6 319 4.8 95.2 1346 219.9 142
VNSRR 14.4 336 5.3 94.7 1338 1.2 139

### v2

@@
-Variable Neighborhood Searching Rerandomization
+Fast Rerandomization via the BRAIN
 Abstract
-Rerandomization discards undesired treatment assignments to ensure covariate balance in randomized experiments.
-However, rerandomization based on acceptance-rejection sampling is computationally inefficient, especially when numerous independent assignments are required to perform randomization-based statistical inference.
-Existing acceleration methods are suboptimal and are not applicable in structured experiments, including stratified experiments and experiments with clusters.
-Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to draw balanced assignments in various experiments efficiently.
-We derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR.
-Simulation studies and a real data example indicate that our method maintains the appealing statistical properties of rerandomization and can sample thousands of treatment assignments within seconds, even in cases where existing methods require an hour to complete the task.
-Keywords causal inference  $\cdot$ clinical trials  $\cdot$ experimental design  $\cdot$ combinatorial optimization  $\cdot$ binary quadratic programming  $\cdot$ variable neighborhood search
+Randomized experiments are a crucial tool for causal inference in many different fields.
+Rerandomization addresses any covariate imbalance in such experiments by resampling treatment assignments until certain balance criteria are satisfied.
+However, rerandomization based on naïve acceptance-rejection sampling is computationally inefficient, especially when numerous independent assignments are required to perform randomization-based statistical inference.
+Existing acceleration methods are suboptimal and not applicable in structured experiments, including stratified and clustered experiments.
+Based on metaheuristics in integer programming, we propose BRAIN — a novel computationally-lightweight methodology that ensures covariate balance in randomized experiments while significantly accelerating the computation.
+Our BRAIN method provides unbiased treatment effect estimators with reduced variance compared to complete randomization, preserving the desirable statistical properties of traditional rerandomization.
+Simulation studies and a real data example demonstrate the benefits of our method in fast sampling while retaining the appealing statistical guarantees.
 1 Introduction
-Randomized experiments balance covariates on average and have been the gold standard for drawing causal inferences.
-However, as pointed out by Morgan and Rubin (2012), covariate imbalances between treatment groups often exist in practice due to the randomness in complete randomization(CR).
-An intuitive and practical solution to this issue is randomizing repeatedly until a prespecified covariate balance criterion is achieved.
-Despite the long history and widespread use of rerandomization (Student, 1938; Cox, 1982; Bailey and Rowley, 1987; Maclure et al., 2006; Imai et al., 2008; Bruhn and McKenzie, 2009), its theoretical implications were first studied by Morgan and Rubin (2012) using the Mahalanobis distance only recently.
-Since then, rerandomization has attracted growing research interests, and its theory underpinnings have been established in various scenarios (Morgan and Rubin, 2015; Li et al., 2018; Zhou et al., 2018; Li et al., 2020; Shi et al., 2022; Wang et al., 2023; Lu et al., 2023).
-As the name of rerandomization indicates, researchers typically conduct rerandomization by repeatedly sampling treatment assignments until a balanced one is generated.
-This procedure is then repeated thousands of times since numerous independent qualified assignments are required to perform randomization-based inference.
-Therefore, rerandomization can be time-consuming using the acceptance-rejection sampling rerandomization(ARSRR) strategy.
-Zhu and Liu (2022) proposed a pair-switching rerandomization(PSRR) method to reduce the computational cost of rerandomization.
-However, the authors only considered a fixed number of covariates and a fixed covariate imbalance threshold.
-As we will show in Section 5, PSRR can be slower than ARSRR in certain settings.
-Moreover, both ARSRR and PSRR are computationally inefficient in settings considered by recent works, including diminishing covariate imbalance and high-dimensional covariates (Branson and Shao, 2021; Wang and Li, 2022; Zhang et al., 2023).
-Additionally, although Zhu and Liu (2022) extended PSRR to sequential experiments, PSRR is not applicable in other structured experiments such as stratified experiments or experiments with clusters.
-In this work, we point out that a treatment assignment with a sufficiently small Mahalanobis distance can be achieved by optimizing a quadratic function of linearly constrained binary variables.
-Based on metaheuristics in combinatorial optimization, we propose a novel variable neighborhood searching rerandomization(VNSRR) method to accelerate rerandomization in various randomized experiments.
-Moreover, we derive the unbiasedness and a lower bound for the variance reduction of the treatment effect estimator under VNSRR.
-Our simulation studies and real data analysis illustrate that VNSRR significantly improves the computational efficiency of rerandomization while maintaining desired statistical properties.
-It can sample thousands of treatment assignments within seconds, even in cases where existing methods require hours to complete the task.
-2 Related Work
-2.1 The Neyman-Rubin Potential Outcome Framework
-This study adopts the Neyman-Rubin potential outcome framework (Neyman, 1923; Rubin, 1974).
+Randomized experiments have been the gold standard for drawing causal inferences in a wide range of settings, including clinical trials [33], social sciences [6], and economic research [14].
+The method relies on the assumption that, by randomly assigning units to treatment and control groups, covariates are balanced.
+Therefore, any differences in outcomes between the groups can be attributed causally to the treatment.
+However, covariate imbalances between treatment groups often exist in practice due to the randomness in such complete randomization (CR) [34, 42, 29, 21].
+This is a major issue because it introduces confounding, making it unclear whether observed differences in outcomes are due to the treatment or to pre-existing differences between groups, consequently invalidating any inference made.
+An intuitive solution to this issue is rerandomization – randomizing repeatedly until some prespecified covariate balance criterion is achieved [39, 12, 3, 26, 19, 10].
+The most popular choice for the balance criterion is requiring the Mahalanobis distance between the covariate matrices of the treatment group and the control group to be lower than some threshold $a$.
+Despite the long history and widespread use of rerandomization, its theoretical implications using the Mahalanobis distance was only recently studied by [29].
+Since then, rerandomization has attracted growing research interests, and its theory underpinnings have been established in various scenarios [30, 22, 45, 23, 38, 40, 24, 44, 8].
+Classical rerandomization methods repeatedly sample treatment assignments until a balanced one is generated.
+However, in randomization-based inference, we need to generate a large number of balanced treatment assignments to ensure reliable estimation and inference.
+This often requires performing thousands of rerandomizations to obtain sufficiently many independent balanced assignments.
+As a consequence, rerandomization is extremely time-consuming using the acceptance-rejection sampling rerandomization (ARSRR) strategy [25, 11, 5].
+This pose practical challenges, especially in large-scale or real-time experiments where treatment assignments must be generated quickly.
+High computational cost may limit the number of balanced assignments available for randomization-based inference, and thus discourage practitioners from adopting rerandomization methods altogether.
+Several methods have been proposed to alleviate the computational burden of rerandomization.
+However, these approaches often fall short in practice, particularly in challenging settings involving diminishing covariate imbalance or high-dimensional covariates [7, 41, 43].
+Meanwhile, some alternative approaches — such as deterministic or semi-deterministic algorithms — may yield a single acceptable assignment more quickly.
+Nevertheless, these methods typically lack the ability to generate independent assignments and do not ensure the unbiasedness of the resulting treatment effect estimator [15, 33, 20].
+To address these limitations, [46] introduced the Pair-Switching Rerandomization (PSRR) algorithm to accelerate the search for balanced assignments in both nonsequential and sequential experiments.
+The algorithm starts with a random assignment and gradually improves it by switching between a treated unit and a control unit repeatedly.
+Motivated by the Metropolis-Hastings algorithm [28, 17], PSRR allows moving to less balanced assignments with non-zero probabilities determined by a temperature hyperparameter to prevent being trapped in a local region.
+While the authors establish the unbiasedness and a lower bound on the variance reduction of the treatment effect estimator under PSRR, the PSRR method remains computationally intensive in many scenarios as shown in Section 5.
+Structured experiments are also common in practice.
+The ability to adapt to these special structures are also important for rerandomization algorithms.
+Rerandomization methods have been proposed for sequential experiments where units arrive group by group [45, 4]; stratified experiment where experimenters divide units into strata and conduct randomization within each stratum [16, 40]; and clustered experiments where all units in the same cluster receives the same treatment [24].
+The applicability of PSRR is limited to sequential experiments only.
+Beside CR and the naïve ARSRR, there does not exist any fast and unified framework that works in all these structured settings.
+Therefore, we aim to propose a method that is fast and adaptable to all structured settings.
+In this paper, we novelly consider reformulating the problem from a constrained optimization (rather than sampling) perspective.
+For an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group; we denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators, where $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise.
+For some covariate balance criterion $M(W)$ — for instance — the popular choice of Mahalanobis distance [29], we consider the optimization problem
+minimize $\quad M(W),$
+subject to $\quad W\in\Big{\{}W\in\{0,1\}^{n}:\sum\nolimits_{i}W_{i}=n_{t}\Big{\}}.$ (1)
+Under this formulation, we consider utilizing combinatorial optimization techniques to conduct rerandomization with early termination.
+There are few studies concerning this specific optimization problem, but some related constrained binary quadratic programming problems have received extensive attention, such as the maximum diversity problem and the heaviest k-subgraph problem.
+These problems share the same feasible sets as (1) but have different objective functions.
+We recommend interested readers refer to [27] for a detailed review of the literature.
+The authors presented extensive experiments to reveal that the best methods to solve the maximum diversity problem are local search based metaheuristics, especially variable neighborhood search [9, 1] and tabu search [32, 1].
+Both methods update their solutions by exchanging coordinates of $W$ that are ones with coordinates of $W$ that are zeroes.
+Variable neighborhood searching systematically changes the neighborhood by executing two steps iteratively.
+It descents to find local optimum in local search steps and perturbs to avoid being trapped in local regions in shaking steps.
+Tabu search maintains a list of forbidden moves, namely the tabu list, to avoid looping over recently visited coordinates of $W$.
+Furthermore, [1] proposed to incorporate tabu search in the local search step of variable neighborhood search.
+Following these intuitions, we propose a novel method to accelerate rerandomization — Balanced RAdomization via INteger programming (BRAIN).
+We show that a treatment assignment with a sufficiently small Mahalanobis distance can be achieved by optimizing a quadratic function of linearly constrained binary variables.
+Moreover, we prove that BRAIN, as a member of the rerandomization (RR) family, leads to treatment effect estimators that are unbiased and has lower variances than CR.
+Our simulation studies and real data analysis illustrate that BRAIN significantly improves the computational efficiency of rerandomization while maintaining the desired statistical properties.
+Using our method, the time it takes to obtain thousands of balanced treatment assignments can be reduced to seconds when existing methods require hours.
+2 Problem Setup
+In this section, we formally formulate the rerandomization framework and discuss existing approaches to solve the problem.
+The Neyman-Rubin Framework.
+We adopt the Neyman-Rubin potential outcome framework [31, 35].
 We consider an experiment with $n$ units, among which $n_{t}$ units are assigned to the treatment group and $n_{c}=n-n_{t}$ units are assigned to the control group.
-We denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators.
-$W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise.
-Each unit $i$ has two potential outcomes, $(Y_{i}(1),Y_{i}(0))$, which correspond to the two treatment arms.
-The individual treatment effect for unit $i$ is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$.
-And the average treatment effect is defined as $\tau=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}(1)-Y_{i}(0))$.
+We denote $W=(W_{1},\ldots,W_{n})^{T}$ as the vector of treatment assignment indicators, where $W_{i}=1$ if unit $i$ is assigned to the treatment group and $W_{i}=0$ otherwise.
+We assume that each unit has $p$-dimensional covariates, denoted by $X_{i}=(X_{i1},\ldots,X_{ip})^{T}$.
+We thus denote the covariate matrix as $X=(X_{1},\ldots,X_{n})^{T}$, with corresponding covariance matrix $S_{XX}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{T}$.
+We assume that each unit $i$ has two potential outcomes, $(Y_{i}(1),Y_{i}(0))$, which correspond to the two treatment arms.
+We denote the individual treatment effect for unit $i$ as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and the average treatment effect as $\tau=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}(1)-Y_{i}(0))$.
 The observed outcome for unit $i$ is $Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0)$.
-We adopt the stable unit treatment value assumption(SUTVA) (Rubin, 1980), which states that the treatment levels are well-defined and there is no interference between $(Y_{i}(1),Y_{i}(0))$ and $\{W_{j}:j\neq i\}$.
-We estimate the treatment effects with the difference-in-means estimator, defined as $\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0)$.
-We assume that each unit has $p$ observed covariates, denoted by $X_{i}=(X_{i1},\ldots,X_{ip})^{T}$.
-The covariates of all units are gathered into a matrix $X=(X_{1},\ldots,X_{n})^{T}$.
-The corresponding covariance matrix is denoted by $S_{XX}=\frac{1}{n-1}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{T}$.
-2.2 Rerandomization Using the Mahalanobis Distance
-For a given assignment $W$, the difference in the covariate means of the two treatment groups is
-$\overline{X}_{t}-\overline{X}_{c}$ $=\frac{1}{n_{t}}X^{T}W-\frac{1}{n_{c}}X^{T}(1_{n}-W)$
-$=\frac{n}{n_{t}n_{c}}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
-And the Mahalanobis distance between the covariates means is defined as
-$M(W)$ $:=\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}\left[Cov[\left(\overline{X}_{t}-\overline{X}_{c}\right)]\right]^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
-$=\frac{n_{t}n_{c}}{n}\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}S_{XX}^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
-$=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n})$
-$=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
+We adopt the stable unit treatment value assumption (SUTVA) [36], which states that the treatment levels are well-defined, and there is no interference between the potential outcomes $(Y_{i}(1),Y_{i}(0))$ and the treatment assignments of other units $\{W_{j}:j\neq i\}$.
+We estimate the treatment effects with the difference-in-means estimator, defined as
+$\widehat{\tau}(W)=\frac{1}{n_{t}}\sum_{i:W_{i}=1}Y_{i}(1)-\frac{1}{n_{c}}\sum_{i:W_{i}=0}Y_{i}(0).$
+We employ randomization-based inference to construct confidence intervals for $\tau$ and test for the sharp null hypothesis $H_{0}:\tau_{i}=0$, $i=1,\ldots,n$.
+See Appendix C for an introduction to randomization-based inference, including the Fisher randomization test.
+The procedure requires generating sufficiently many (e.g., 1000) i.i.d.
+assignments.
+Rerandomization with the Mahalanobis Distance.
+[29] suggested to conduct rerandomization by sampling $W$ from $\{W:\sum_{i=1}^{n}W_{i}=n_{t},W_{i}\in\{0,1\}\}$ repeatedly until $M(W)\leq a$, where $a$ is a prespecified threshold and $M(W)$ is the Mahalanobis distance between the covariates matrices of the treatment groups.
+$M(W)$ is formally defined as
+$M(W)$ $:=\left(\overline{X}_{t}-\overline{X}_{c}\right)^{T}\left[\textnormal{Cov}[\left(\overline{X}_{t}-\overline{X}_{c}\right)]\right]^{-1}\left(\overline{X}_{t}-\overline{X}_{c}\right)$
+$=\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}H\left(W-\frac{n_{t}}{n}1_{n}\right),$
 where $H=\frac{n}{n_{t}n_{c}}XS_{XX}^{-1}X^{T}$.
-Morgan and Rubin (2012) suggested to conduct rerandomization by sampling $W$ from $\{W:\sum_{i=1}^{n}W_{i}=n_{t},W_{i}\in\{0,1\}\}$ repeatedly until $M(W)\leq a$, where $a$ is a prespecified threshold.
-We refer to rerandomization based on this acceptance-rejection sampling strategy as ARSRR.
-2.3 Pair-switching Rerandomization (PSRR)
-To avoid the computationally inefficient acceptance-rejection sampling approach and accelerate rerandomization in both nonsequential and sequential experiments, Zhu and Liu (2022) suggested to search for balanced assignments with a pair-switching strategy, i.e., PSRR.
-The algorithm essentially starts with a random assignment and gradually improves it by switching between a treated unit and a control unit repeatedly.
-Motivated by the Metropolis-Hastings algorithm, PSRR allows moving to less balanced assignments with a non-zero probability to prevent being trapped in a local region.
-The authors obtained the unbiasedness and a lower bound on the variance reduction of the treatment effect estimator under PSRR.
-However, as pointed out in Section 1 and demonstrated by the simulation studies in Section 5, PSRR is still time-consuming in many settings and is not applicable in certain structured experiments.
-3 Variable Neighborhood Searching Rerandomization(VNSRR)
-We introduce our variable neighborhood searching rerandomization (VNSRR) in this section.
-3.1 The VNSRR Algorithm
+Though this approach is commonly used in practice, it is computationally expensive.
+We will now introduce our BRAIN method, which addresses the computational challenge by leveraging metaheuristics from integer programming.
+3 The BRAIN Method
+In this section, we first introduce our Balanced RAdomization via INteger programming (BRAIN) algorithm for rerandomization.
+We then establish statistical properties of the resulting difference-in-means estimator using BRAIN.
+3.1 The BRAIN Algorithm
 We notice that rerandomization is closely related with the following constrained binary quadratic programming problem.
-minimize $\quad M(W)=(W-\frac{n_{t}}{n}1_{n})^{T}H(W-\frac{n_{t}}{n}1_{n}),$
+minimize $\quad M(W)=\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}H\left(W-\frac{n_{t}}{n}1_{n}\right),$
 subject to $\quad\sum_{i=1}^{n}W_{i}=n_{t},$
@@
 For simplification, we denote the feasible set of (3.1) as $\mathcal{F}_{0}=\{W\in\{0,1\}^{n}:\sum_{i}W_{i}=n_{t}\}$.
-While rerandomization aims to find a feasible $W$ in $\mathcal{F}_{0}$ with sufficiently small $M(W)$ rather than the minimizer of $M(W)$ in $\mathcal{F}_{0}$, we can still utilize combinatorial optimization techniques to conduct rerandomization and are allowed to terminate the optimization process as long as $M(W)$ falls below a prespecified threshold $a$.
-There are few studies concerning the specific optimization problem in (3.1), but some related constrained binary quadratic programming problems have received extensive attention, such as the maximum diversity problem and the heaviest k-subgraph problem.
-These problems share the same feasible sets as (3.1) but have different objective functions.
-We recommend that interested readers refer to Martí et al.
-(2013) for a detailed review of the literature.
-The authors presented extensive experiments to reveal that the best methods to solve the maximum diversity problem are local search based metaheuristics, especially variable neighborhood search (Brimberg et al., 2009; Aringhieri and Cordone, 2011) and tabu search (Palubeckis, 2007; Aringhieri and Cordone, 2011).
-Both methods update their solutions by exchanging coordinates of $W$ that are ones with coordinates of $W$ that are zeroes.
-Variable neighborhood search systematically changes the neighborhood by executing two steps iteratively.
-It descents to find local optimum in local search steps and perturbs to avoid being trapped in local regions in shaking steps.
-Tabu search maintains a list of forbidden moves, namely the tabu list, to avoid looping over recently visited coordinates of $W$.
-Aringhieri and Cordone (2011) proposed to incorporate tabu search in the local search step of variable neighborhood search.
+While rerandomization aims to find a feasible $W$ in $\mathcal{F}_{0}$ with sufficiently small $M(W)$ rather than the minimizer of $M(W)$ in $\mathcal{F}_{0}$, we can still utilize combinatorial optimization techniques to conduct rerandomization with early termination (as long as $M(W)$ falls below a prespecified threshold $a$).
+Motivated by the empirical success of local-search-based metaheuristics in related optimization problems, we propose the BRAIN method as follows.
+BRAIN starts with a random treatment assignment.
+Then, it conducts a local search step where we sequentially examine $L$ disjoint random pairs of treated units and control units and swaps a pair if it leads to a more balanced assignment.
+If any swap has been made in the local search step, BRAIN will continue to search in the current local region.
+When no pairs are switched in the local search step, indicating a likely local optimum, BRAIN performs a perturbation step by randomly switching $S$ pairs of treated and control units.
+The algorithm will repeat the local search step and the perturbation step until reaching a satisfiable sample with $M(W)\leq\alpha$.
+The full algorithm is shown in Algorithm 1.
+BRAIN conducts the local search step differently from any existing variable neighborhood searching methods.
+We borrow the idea of preventing looping over recently visited units from tabu search and examine $L$ disjoint pairs.
+However, BRAIN differs from tabu search in the sense that BRAIN avoids maintaining the tabu list and checking whether each swap is forbidden, further improving computational efficiency.
+$L$ and $S$ are two hyperparameters in BRAIN.
+If $L$ is too small, the algorithm would still loop over recently visited pairs.
+If $S$ is too large, the perturbation is too strong such that it is similar to random restart.
+We compare different choices of $L$ and $S$ in Appendix D and suggest setting $(L,S)=(n/2,1)$.
+Our method differs from PSRR in that it accepts swaps only if $M(W)$ decreases in local search step.
+This way, it avoids inefficiently moving to a treatment assignment $W$ with higher $M(W)$, which frequently happens for PSRR.
+By avoiding this issue, our algorithm leads to significant faster descent of $M(W)$ and therefore much shorter computation time.
+PSRR with zero temperature is equivalent to BRAIN with $L=1$ and without the perturbation steps.
+Our experiments show that this special version of PSRR is faster than its default version but still much slower than BRAIN.
 Input: Covariates $X$, group sizes $(n_{t},n_{c})$, threshold $a$, hyperparameter $(L,S)$.
-Set $t=0$;
-Set $W^{(0)}$ as $n_{t}$ elements equal to 1 and $n_{c}$ elements equal to 0 with random positions;
-Set $M^{(0)}=M(W^{(0)})$;
+Set $t=0$, $W^{(0)}$ randomly with $n_{t}$ many 1’s and $n_{c}$ many 0’s, $M^{(0)}=M(W^{(0)})$;
 while $M^{(t)}>a$ do
@@
              if $M^{*}<M^{(t)}$ then
-                   Set $\widetilde{S}=0$;
-                   Set $t=t+1$;
-                   Set $W^{(t)}=W^{*}$;
-                   Set $M^{(t)}=M^{*}$;
-                   if $M^{(t)}<a$ then break;
+                   Set $\widetilde{S}=0$, $t=t+1$, $W^{(t)}=W^{*}$, $M^{(t)}=M^{*}$;
+                   if $M^{(t)}\leq a$ then break;
+             end if
+       end for
       if $\widetilde{S}>0$ then
@@
                    Set $t=t+1$;
+             end for
+       end if
+end while
 Output: $W=W^{(t)}$.
-Algorithm 1 Variable neighborhood searching rerandomization
-Motivated by the empirical success of local-search-based metaheuristics in related optimization problems, we propose a novel VNSRR method for rerandomization through variable neighborhood searching, as summarized by Algorithm 1.
-VNSRR starts with a random assignment and then executes the local search step and the shaking step iteratively until an acceptable assignment is reached.
-VNSRR conducts the local search step differently from any existing variable neighborhood search methods.
-It is inspired by but different from tabu search methods.
-In the local search step, VNSRR sequentially examines $L$ disjoint random pairs of treated units and control units and swaps a pair if it leads to a more balanced assignment.
-We borrow the idea of preventing looping over recently visited units from tabu search.
-However, VNSRR differs from tabu search in that VNSRR avoids maintaining the tabu list and checking whether each swap is forbidden, further improving computational efficiency.
-We note that when selecting multiple treated units or control units in the local search step, the ordering of the units is also randomly selected.
-In other words, we randomly select a permutation of $L$ positions of the 1’s or 0’s in $W^{(t)}$.
-In the shaking step, VNSRR generates an assignment in the neighborhood of the current assignment by randomly switching $\widetilde{S}$ pairs of treated and control units.
-When no pairs are switched in the local search step, indicating a likely local optimum, $\widetilde{S}$ is set as a positive value $S$ to prevent getting stuck.
-Otherwise, $\widetilde{S}$ is set as 0, and no perturbation is performed to allow for further search in the current local region.
-We compare different choices of $L$ and $S$ in our experiments, and the results are relegated to the Supplement.
-We point out that both PSRR and VNSRR use pairwise swap as the basic operation in their algorithms, but use it with different strategies.
-PSRR allows switching recently considered units, while VNSRR does not.
-PSRR, as motivated by simulated annealing and Metropolis-Hastings algorithm, accepts swaps that lead to worse covariate balance, i.e., larger $M(W)$, with a certain probability determined by a temperature parameter.
-In contrast, VNSRR keeps moving to new $W$ with smaller $M(W)$ in the local search step until none of the $L$ swaps lead to smaller $M(W)$, and then the shaking step is conducted.
-Our experiments illustrate that PSRR can move to $W$ with worse $M(W)$ quite frequently, while the strategy of VNSRR leads to a faster descent of $M(W)$.
-In fact, PSRR with zero temperature is equivalent to VNSRR with $L=1$ and without shaking steps.
-Our experiments show that this special version lies in the middle of PSRR and VNSRR.
-The zero-temperature version of PSRR was not considered by the authors of PSRR and is faster than their default version.
-Our default VNSRR is faster than VNSRR with $L=1$ and without shaking steps.
-Like PSRR, we adopt the following trick to update the Mahalanobis distance efficiently.
-When we obtain a new assignment $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$, $M(W^{*})$ can be efficiently calculated as
-$M(W^{*})=$ $M^{(t)}-2\sum_{i=1}^{n}W_{i}^{(t)}H_{O_{l}i}+H_{O_{l}O_{l}}+2\sum_{i=1}^{n}W_{i}^{*}H_{Z_{l}i}-H_{Z_{l}Z_{l}}+h_{O_{l}}-h_{Z_{l}},$
-where $h=2\frac{n_{t}}{n}H1_{n}$.
-We note that some deterministic or semi-deterministic algorithms might generate a single acceptable assignment faster than VNSRR.
-However, such algorithms cannot be used to generate numerous independent assignments and do not guarantee the unbiasedness of the corresponding treatment effect estimator.
-The next section discusses the statistical properties of different rerandomization methods.
-3.2 Statistical Properties of VNSRR
-In this section, we establish the statistical properties of various rerandomization algorithms based on the Mahalanobis distance.
-For clarification, a rerandomization algorithm here is a sampling algorithm that randomly generates a balanced treatment assignment from the feasible set $\mathcal{F}\subseteq\mathcal{F}_{0}$ by generating a chain of treatment assignments, $(W^{(0)},\ldots,W^{(T)})$, where $W^{(t)}\in\mathcal{F}$ for all $t$ and $W^{(T)}\in\mathcal{F}\cap\{W:M(W)\leq a\}$ is the final output.
-We provide sufficient conditions for a rerandomization algorithm to yield unbiased treatment effect estimators with bounded variance.
-Specifically, we consider the following conditions, which assumes Morgan and Rubin (2012) as well as Zhu and Liu (2022).
-Condition 1
+Algorithm 1 Balanced RAdomization via INteger programming (BRAIN)
+3.2 Statistical Properties of BRAIN
+In this section, we establish the statistical properties of BRAIN. We consider the following conditions, which are standard in the literature [29, 46].
+Condition 1.
 (linear decomposition) For $w_{i}=0,1$, $Y_{i}(w_{i})=\beta_{0}+\beta^{T}X_{i}+\tau w_{i}+e_{i}$, where $\beta_{0}+\beta^{T}X_{i}$ is the linear projection of $Y_{i}(0)$ onto $(1,X)$ and $e_{i}$ is the deviation from the linear projection.
-Condition 2
+Condition 2.
 (normality) $\widehat{\tau}$ and $\overline{X}_{t}-\overline{X}_{c}$ are normally distributed.
-And we focus on three key properties of rerandomization algorithms.
-Property 1
-(symmetric feasible set) $\mathcal{F}=1-\mathcal{F}$.
-Property 2
-(uniform initialization) For any $w\in\mathcal{F}$,
-$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
-Property 3
-(updating rule) There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
-$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
-Based on Property 1-3, we prove that the difference-in-means estimator is unbiased.
-Theorem 1
-If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
-Based on Condition 1-2 and Property 1-3, we prove that the difference-in-means estimator has smaller variance than that under complete randomization and construct a lower bound for the variance reduction.
-Denote $V_{CR}$ as the variance of the difference-in-means estimator under complete randomization, we prove the following theorem.
-Theorem 2
-Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
-$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
-where $R^{2}=\frac{\beta^{T}Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]\beta}{V_{CR}}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$.
-Similar to the definition of $V_{CR}$, $Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]$ represents the covariance matrix of $\overline{X}_{t}-\overline{X}_{c}$ under complete randomization.
-One can verify that VNSRR with $\mathcal{F}_{0}$ as the feasible set satisfies our proposed requirements under the following conditions.
-Condition 3
+Condition 3.
 (equal group sizes) $n_{t}=n_{c}=\frac{n}{2}$.
-Proposition 1
-VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
-Therefore, the unbiasedness and variance reduction of the difference-in-means estimator under VNSRR are immediately evident as corollaries of Theorem 1 and Theorem 2.
-Corollary 1
-Under Condition 3 and VNSRR, $E[\widehat{\tau}]=\tau$.
-Corollary 2
-Under Condition 1, Condition 2, Condition 3, and VNSRR,
-$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
-One can verify that CR, ARSRR, and PSRR all satisfy Property 1, Property 2, and Property 3 under certain conditions.
-As a result, their statistical properties are established by this framework as well.
-All verifications and proofs are included in the Supplement.
-4 VNSRR for Structured Randomized Experiments
-In this section, we extend VNSRR to randomized experiments with various types of structures.
-For each VNSRR variant, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator.
-We will reuse several notations across different settings, leveraging the similarities between settings to streamline our notation system.
-The pseudocode of each variant, along with the theoretical results and their proofs, is provided in the Supplement.
-4.1 VNSRR for Sequentially Randomized Experiments
+We prove that the difference-in-means estimator is unbiased under BRAIN.
+Theorem 1.
+Under Condition 3, if $W$ is generated by BRAIN, then $\mathbb{E}[\widehat{\tau}]=\tau$.
+We also prove that the difference-in-means estimator has smaller variance than that under complete randomization and construct a lower bound for the variance reduction.
+Denote $\textnormal{V}_{\textnormal{CR}}$ as the variance of the difference-in-means estimator under complete randomization, we prove the following theorem.
+Theorem 2.
+Under Conditions 1-3, if $X$ has $p$ covariates and $W$ is generated by BRAIN with threshold $a$, then
+$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2},$
+where
+$R^{2}=\frac{\beta^{T}\textnormal{Cov}_{\textnormal{CR}}[\overline{X}_{t}-\overline{X}_{c}]\beta}{\textnormal{V}_{\textnormal{CR}}}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}\textnormal{V}_{\textnormal{CR}}}.$
+Here $\textnormal{Cov}_{\textnormal{CR}}[\overline{X}_{t}-\overline{X}_{c}]$ represents the covariance matrix of $\overline{X}_{t}-\overline{X}_{c}$ under complete randomization.
+Like other rerandomization method, the difference-in-mean estimator still satisfies the same unbiasedness and variance reduction properties.
+This means that BRAIN is able to accelerate the process of rerandomization without sacrificing the desirable statistical properties.
+4 BRAIN for Structured Randomized Experiments
+In this section, we extend BRAIN to randomized experiments with various types of structures – sequential, stratified, and cluster randomized experiments.
+We will discuss how to adapt our algorithm to each of the three settings.
+The full algorithms are provided in Appendix A, as well as the validation of their unbiasedness and variance reduction properties.
+4.1 BRAIN for Sequentially Randomized Experiments
 In Section 3, we consider offline rerandomization, which assigns treatment to all units at once.
-However, as pointed out by Zhou et al.
-(2018), experimenters may be unable to wait for all units to arrive before performing the experiments.
-It is common to assign treatment to units group by group in practice.
-To balance the covariates in such sequential experiments, Zhou et al.
-(2018) suggested performing sequential rerandomization, which can be computationally challenging.
-Bertsimas et al.
-(2019) proposed a novel algorithm for online allocation that leverages robust mixed-integer optimization.
-However, the authors focused on optimal designs, which are inherently different from the rerandomization framework.
-We now review the basic concepts and notations of sequential rerandomization.
-We assume that the experiment consists of $K$ stages, and the $n$ units are divided into $K$ sequential groups of sizes $(n_{1},\ldots,n_{K})$.
-In the $k$-th stage, $n_{tk}$ units in the $k$-th group are assigned to the treatment arm, and the remaining $n_{ck}$ units are assigned to the control arm.
-We use the subscript $[k]$ to represent the reduced version of a term calculated based on only the first $k$ groups.
-The covariates of the first $k$ groups are denoted by $X_{[k]}$ and its covariance matrix is denoted by $S_{XX[k]}$.
-The assignment status of the first $k$ groups is denoted by $W_{[k]}$.
-And we define $n_{[k]}=\sum_{g=1}^{k}n_{g}$, $n_{t[k]}=\sum_{g=1}^{k}n_{tg}$, $n_{c[k]}=\sum_{g=1}^{k}n_{cg}$, $H_{[k]}=\frac{n_{[k]}}{n_{t[k]}n_{c[k]}}X_{[k]}S_{XX[k]}^{-1}X_{[k]}^{T}$.
-The Mahalanobis distance corresponding to $W_{[k]}$ is defined as
-$M_{[k]}$ $=(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})^{T}H_{[k]}(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})$
-Sequential rerandomization proceeds as follows: When the $k$-th group enrolls in the experiment, the experimenter randomly assigns treatment to the units without changing the assignment status of the previously enrolled groups.
-The experimenter aims to balance the covariates of all units in the first $k$ groups.
-Therefore, the experimenter rerandomizes the $n_{k}$ units until $M_{[k]}\leq a_{k}$, where $a_{k}$ is a prespecified threshold.
-Zhou et al.
-(2018) proposed to use acceptance-rejection sampling in this procedure, so we refer to their method as SeqARSRR.
-Recently, Zhu and Liu (2022) proposed a sequential pair-switching rerandomization(SeqPSRR) method to reduce the computational cost.
-SeqPSRR essentially performs PSRR in each stage of the sequential experiment.
-In Section 5.2, we will show that both SeqARSRR and SeqPSRR can be time-consuming.
-We note that sequential rerandomization in the $k$-th stage can be formulated into the following constrained binary quadratic programming problem, whose feasible set we refer to as $\mathcal{F}_{[k]}$.
-minimize $\quad M_{[k]}(W_{[k]}),$
-subject to $\quad\sum_{i=n_{[k-1]}+1}^{n_{[k]}}W_{[k],i}=n_{tk},$
-$\quad W_{[k],i}\in\{0,1\},\;i=n_{[k-1]}+1,\ldots,n_{[k]},$
-$\quad W_{[k],i}=W_{[k-1],i},\;i=1,\ldots,n_{[k-1]}.$ (2)
-Based on this observation, we extend VNSRR to sequentially randomized experiments and refer to this variant as SeqVNSRR.
-SeqVNSRR performs VNSRR in each stage of the experiment while fixing the assignments obtained in previous stages.
-It also adopts the updating trick, the specific form of the local search step, and the hyperparameter combination mentioned in Section 3 to reduce the computational cost.
-To analyze the statistical properties of SeqVNSRR, we consider the following variants of Condition 3 and Property 1-3 in sequential experiments.
-Condition 4
-$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
-Property 4
-For any $k$, $\mathcal{F}_{[k]}=1-\mathcal{F}_{[k]}$.
-Property 5
-For any $k$ and $w\in\mathcal{F}_{[k]}$,
-$P\left(W_{[k]}^{(0)}=w\right)=\frac{1}{|\mathcal{F}_{[k]}|}.$
-Property 6
-There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $k$, $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain for the $k$-th stage,
-$P\left(W_{[k]}^{(t)}=w^{(t)}\mid W_{[k]}^{(0)}=w^{(0)},\ldots,W_{[k]}^{(t-1)}=w^{(t-1)}\right)=u_{t}(M_{[k]}(w^{(0)}),\ldots,M_{[k]}(w^{(t)})),$
-Based on these conditions and properties, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under SeqVNSRR.
-4.2 VNSRR for Stratified Randomized Experiments
-Other than rerandomization, stratification stands out as another approach to balance covariates.
-The idea of stratification, or blocking, was first introduced by Fisher (1926) to balance discrete covariates.
+However, in sequential experimenters where participants arrive one by one or in groups, it is common to assign treatment to units group by group in practice.
+We assume that the experiment consists of $K$ stages such that the $n$ units are divided into $K$ sequential groups of sizes $(n_{1},\ldots,n_{K})$.
+In the $k$-th stage, $n_{t_{k}}$ units in the $k$-th group are assigned to the treatment arm, and the remaining $n_{c_{k}}$ units are assigned to the control arm.
+We use the subscript $[k]$ to represent the reduced version of any quantity calculated based on only the first $k$ groups.
+For example, $M_{[k]}(W_{[k]})$ represents the Mahalanobis distance of the covariate matrices of the first $k$ group.
+Sequential rerandomization aims to do the following: when the k-th group enrolls in the experiment, randomly assign treatment to the units without changing the assignment status of the previously enrolled groups, while balancing the covariates of all units in the first k groups with $M_{[k]}(W_{[k]})\leq a_{k}$.
+We extend our BRAIN method to sequentially randomized experiments (denoted by SeqBRAIN, details in Algorithm 2) where it performs BRAIN in each stage of the experiment while fixing the assignments obtained in previous stages.
+4.2 BRAIN for Stratified Randomized Experiments
+Other than rerandomization, stratification is a common alternative approach to balance covariates.
 In stratified randomized experiments, experimenters divide units into strata, namely, layers or levels, according to the discrete covariates and conduct complete randomization within each stratum.
-Wang et al.
-(2023) proposed two strategies that combine stratification and rerandomization to improve covariate balance further.
-The overall strategy rerandomizes over all strata and measures the overall covariate imbalance, while the stratum-specific strategy rerandomizes within each strata separately and measures the stratum-specific covariate imbalance.
-Both strategies have their unique advantages.
-However, since the stratum-specific strategy is identical to the setting we consider in Section 3 within each stratum, we mainly focus on the overall strategy in this section.
+If stratum-specific covariate balance is of interest, we can directly apply BRAIN to each strata.
+If the overall covariate balance is of interest, we can modify BRAIN to effectively account for this structure.
 We assume that the $n$ units are divided into $K$ strata of sizes $(n_{1},\ldots,n_{K})$.
-The index set of units in the $k$-th stratum is denoted by $\mathcal{I}_{k}$.
-In the $k$-th stratum, $n_{tk}$ units are assigned to the treatment arm and $n_{ck}=n_{k}-n_{tk}$ units are assigned to the control arm.
+The set of indices for units in the $k$-th stratum is denoted by $\mathcal{I}_{k}$.
+In the $k$-th stratum, $n_{t_{k}}$ units are assigned to the treatment arm and $n_{c_{k}}=n_{k}-n_{t_{k}}$ units are assigned to the control arm.
 Similar to Section 2, the assignment status is represented by a $n$-dimensional vector $W$ and the covariates are gathered into a $n\times p$ matrix $X$.
-The overall strategy proposed by Wang et al.
-(2023) samples assignments from the space $\{W\in\{0,1\}^{n}:\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\forall k\}$ repeatedly until $M(W)<a$.
+The ’overall’ strategy proposed by [40] samples assignments from the space $\mathcal{W}_{\textnormal{strat}}=\{W\in\{0,1\}^{n}:\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{t_{k}},\forall k\in[K]\}$ repeatedly until $M(W)\leq a$.
 To our knowledge, the only existing method to achieve this goal is to perform acceptance-rejection sampling, which we refer to as StratARSRR.
-We note that rerandomization in stratified experiments can be formulated into the following constrained binary quadratic programming problem.
-minimize $\quad M(W),$
-subject to $\quad\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\;k=1,\ldots,K$
-$\quad W_{i}\in\{0,1\},\;i=1,\ldots,n.$ (3)
-Based on this observation, we extend VNSRR to stratified randomized experiments and refer to this variant as StratVNSRR.
-StratVNSRR performs variable neighborhood searching while considering the constraints imposed by the stratum structure.
-In the local search step, StratVNSRR randomly selects a list of multiple disjoint pairs of treated units and control units from each stratum.
-Then StratVNSRR aggregates the lists from different strata and permutes the pairs in the aggregated list.
-StratVNSRR follows this new ordering of unit pairs to examine whether switching each pair leads to a more balanced assignment.
+We note that rerandomization in stratified experiments is equivalent to minimizing $M(W)$ under the constraint that $W\in\mathcal{W}_{\textnormal{strat}}$.
+We can extend BRAIN to stratified randomized experiments and refer to this variant as StratBRAIN (Algorithm 3).
+StratBRAIN performs local searching and perturbation iteratively while considering the constraints imposed by the stratum structure.
+In the local search step, StratBRAIN randomly selects a list of multiple disjoint pairs of treated units and control units from each stratum.
+Then StratBRAIN aggregates the lists from different strata and permutes the pairs in the aggregated list.
+StratBRAIN follows this new ordering of unit pairs to examine whether switching each pair leads to a more balanced assignment.
 Compared to consecutively examining pairs in the same stratum, this procedure avoids any predetermined ordering of units or strata and allows for more flexible search in the local region.
-In the shaking step, StratVNSRR conducts random unit swaps within every stratum.
+In the perturbation step, StratBRAIN conducts random unit swaps within every stratum.
 This ensures that the new assignment is sufficiently distant from the current one.
-We note that the feasible set in (4.2) is a subset of the feasible set in (3.1).
-As a result, StratVNSRR belongs to the class of rerandomization methods we consider in Section 3.
-We verify that StratVNSRR always satisfies Property 2-3 and satisfies Property 1 under Condition 4.
-Therefore, the statistical properties of the difference-in-means estimator under StratVNSRR can be derived as corollaries of Theorem 1 and Theorem 2.
-4.3 VNSRR for Cluster Randomized Experiments
-Many public health and social science experiments (Donner et al., 2000; Hayes and Moulton, 2017; Schochet, 2020) assign the treatment at the cluster level rather than the individual level due to practical limitations or policy factors.
+4.3 BRAIN for Cluster Randomized Experiments
+Many public health and social science experiments [13, 18, 37] assign the treatment at the cluster level rather than the individual level due to practical limitations or policy factors.
 Suppose the $n$ units are divided into $K$ clusters of sizes $(n_{1},\ldots,n_{K})$ in a cluster randomized experiment.
 We denote $\mathcal{I}_{k}$ as the index set of units in the $k$-th cluster.
-The experimenter assigns $K_{t}$ clusters to the treatment arm and $K_{c}=K-K_{t}$ clusters to the control arm.
+Cluster randomization assigns $K_{t}$ clusters to the treatment arm and $K_{c}=K-K_{t}$ clusters to the control arm.
 We note that when the $n_{i}$’s are not all equal, neither the treatment arm nor the control arm has a fixed number of units, which is different from all the settings we consider in previous sections.
@@
 $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit.
-Lu et al.
-(2023) suggested conducting cluster rerandomization to balance covariates in cluster randomized experiments.
+[24] suggested conducting cluster rerandomization to balance covariates in cluster randomized experiments.
 The authors proposed two metrics that measure the covariate imbalance.
 One is the Mahalanobis distance of cluster-level covariates, and the other is the Mahalanobis distance of individual-level covariates.
-Since using the first metric is equal to treating a cluster as an aggregated unit and can be efficiently solved by the approach proposed in Section 3, we mainly focus on the second metric in this section.
-We denote the individual-level covariates as a $n\times p$ matrix, $X$.
-Some cluster-level covariates may also be available, but we can construct individual-level covariates from cluster-level ones by ensuring they are identical within each cluster.
-We then formulate cluster rerandomization into the following constrained binary quadratic programming problem.
-minimize $\quad M(W),$
-subject to $\quad W_{i}=U_{k},\;\text{for all }k\text{ and }i\in\mathcal{I}_{k}$
-$\quad\sum_{k=1}^{K}U_{k}=K_{t},$
-$\quad U_{k}\in\{0,1\},\;k=1,\ldots,K.$ (4)
-We propose the ClustVNSRR method, a variant of VNSRR, to solve (4.3) efficiently and accelerate cluster rerandomization.
-ClustVNSRR performs variable neighborhood searching at the cluster level and treats the clusters as the operational units in the local and shaking search steps.
-Similar to Section 4.2, we note that the feasible set in (4.3) is a subset of the feasible set in (3.1).
-Therefore, we derive the statistical properties of the difference-in-means estimator under ClustVNSRR by verifying that ClustVNSRR always satisfies Property 2-3 and satisfies Property 1 under the following variant of Condition 3.
-Condition 5
-$K_{t}=K_{c}=\frac{K}{2}$.
+Since using the first metric is equal to treating a cluster as an aggregated unit and can be efficiently solved by the approach proposed in Section 3, we focus on how to modify BRAIN to minimize the second metric.
+We thus propose the ClustBRAIN method (Algorithm 4), a variant of BRAIN, which performs local searching and perturbation at the cluster level while treating the clusters as the operational units.
+The detailed algorithms for each of the three cases are deferred to Appendix A.
+In all three structured settings, the difference-in-means estimator resulting from our proposed methods enjoys similar unbiasedness and variance reduction properties to those in Section 3.
+The details and proofs are shown in Appendix B.
 5 Simulation Studies and Real Applications
-This section compares VNSRR with CR, ARSRR, and PSRR.
-To compare the computation time of VNSRR and the existing methods, we use each method to sample 1000 acceptable assignments respectively in one simulation and replicate the simulation 100 times on a MacBook Pro (Apple M1 Max Chip, 32 GB Memory, 10 Cores).
-To examine the statistical properties, we use each method to sample one acceptable assignment in the design stage to conduct the experiment and 1000 acceptable assignments in the analysis stage to conduct randomization-based inference.
+This section compares BRAIN with CR, ARSRR, and PSRR.
+To compare the computation time of BRAIN and the existing methods, we use each method to sample 1000 acceptable assignments respectively in one simulation and replicate the simulation 100 times on a MacBook Pro (Apple M1 Max Chip, 32 GB Memory, 10 Cores).
+To examine the statistical properties (such as standard deviation of the treatment effect estimator and length of confidence interval), we use each method to sample one acceptable assignment in the design stage to conduct the experiment and 1000 acceptable assignments in the analysis stage to conduct randomization-based inference.
 We replicate the simulation 1000 times using a High-Performance Computing cluster.
+We note that CR is expected to be the fastest method with the worst statistical performances as it has no requirements for covariate balance at all.
+ARSRR, PSRR, and BRAIN are expected to have similar performances in terms of the statistical properties because all of them belong to the rerandomization family and use the same criterion $M(W)\leq a$.
+Therefore, the ideal performance of BRAIN would be:
+  1. 1.
+    Statistical properties: similar to ARSRR and PSRR, better than CR;
+  2. 2.
+    Computational time: (much) faster than ARSRR and PSRR, but not CR.
 5.1 Simple Randomized Experiments
-We first consider the simple randomized experiments described in Section 3 and mimic the simulation settings in Zhu and Liu (2022).
+We first consider the simple randomized experiments described in Section 3 and mimic the simulation settings in [46].
 We assume that the covariates follow the standard normal distribution identically and independently, $X_{ij}\overset{i.i.d.}{\sim}N(0,1)$, $i=1,\ldots,n$, $j=1,\ldots,p$.
 The potential outcomes of the control group are generated independently from a linear model, $Y_{i}(0)=\sum_{j=1}^{p}X_{ij}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,p)$.
-We conduct simulations under both the null hypothesis and the alternative hypothesis, where we set $Y_{i}(1)=Y_{i}(0)$ and $Y_{i}(1)=Y_{i}(0)+0.2\sqrt{Var[Y_{i}(0)]}$, respectively.
-We consider three sample sizes, $n\in\{30,100,500\}$.
-The size of the treatment group and the control group are set as equal.
-Morgan and Rubin (2012) proved that the asymptotic acceptance probability of the assignment is $p_{a}=P(\chi_{p}^{2}<a)$ and Li et al.
-(2018) recommended that $p_{a}=10^{-3}$.
-Wang and Li (2022) considered settings where $p_{a}$ and $p$ are allowed to grow with $n$ rather than fixed.
-Motivated by these works, we considered two thresholds of the Mahalanobis distance, $p_{a}\in\{10^{-3},10^{-4}\}$, and two regimes for the number of covariates, $p=2$ and $p=n/2$.
+We conduct simulations under both the null hypothesis and the alternative hypothesis:
+$H_{0}:Y_{i}(1)=Y_{i}(0),$
+$H_{a}:Y_{i}(1)=Y_{i}(0)+0.2\sqrt{\textnormal{Var}[Y_{i}(0)]}.$
+We consider three sample sizes, $n\in\{30,100,500\}$. The size of the treatment group and the control group are set as equal.
+[29] proved that the asymptotic acceptance probability of the assignment is $p_{a}=P(\chi_{p}^{2}\leq a)$ and [22] recommended setting $a$ such that $p_{a}=10^{-3}$.
+[41] considered settings where $p_{a}$ and $p$ are allowed to grow with $n$ rather than fixed.
+Motivated by these works, we consider two acceptance probabilities, $p_{a}\in\{10^{-3},10^{-4}\}$, and calculate the threshold $a$ by solving $p_{a}=P(\chi_{p}^{2}\leq a)$.
+We consider two regimes for the number of covariates, $p=2$ and $p=n/2$.
 We perform Fisher randomization tests with a significance level $\alpha=0.1$ and construct randomization-based confidence intervals with a nominal coverage rate of 90%.
-The exact formulas of Fisher randomization test p-values and randomization-based confidence intervals are provided in the Supplement.
-We use the default tuning parameter of PSRR, as Zhu and Liu (2022) pointed out that different tuning parameters lead to similar performance.
-For each method, we examine the absolute bias and standard deviation of the difference-in-means estimator, the size and power of the Fisher randomization test, the coverage and length of confidence intervals, and the average run time to sample $10^{3}$ acceptable assignments.
-We also use the randomness metric in Zhu and Liu (2022) to measure the randomness of different methods in the Supplement.
-Table 1 shows the statistical and computational performances of all methods in two typical settings: a low-dimensional setting with $(n,p)=(30,2)$ and a high-dimensional setting with $(n,p)=(500,250)$.
-The complete results are provided in the Supplement.
-VNSRR is significantly faster than ARSRR and PSRR in all settings.
-VNSRR merely requires seconds to generate 1000 acceptable assignments, even in cases where ARSRR or PSRR requires an hour.
+For each method, we examine the absolute bias and standard deviation of the difference-in-means estimator, the size and power of the Fisher randomization test, the coverage and length of confidence intervals, and the average run time to sample 1000 acceptable assignments.
+We also use the randomness metric in [46] to measure the randomness of different methods.
+Recall that we need 1000 i.i.d.
+assignments to conduct randomization-based inference.
+If a method is almost deterministic or lacks randomness, its statistical inference performance can be impaired.
+Table 1: Statistical and computational performance of different methods in simple randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
+$n$ $p$ $p_{a}$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
+Bias SD Size Power CP Len
+30 2 $10^{-3}$ CR 6.2 74 10.5 14 89.5 246 $<0.1$
+ARSRR 0.1 51 8.7 19 91.3 177 12.7
+PSRR 1.5 51 10.1 18 89.9 177 50.3
+BRAIN 1.6 52 11.0 17 89.0 177 1.4
+500 250 $10^{-4}$ CR 0.8 201 10.4 88 89.6 660 $<0.1$
+PSRR 4.3 187 10.3 91 89.7 611 2878.8
+BRAIN 5.7 181 9.8 92 90.2 608 1.3
+Table 1 shows the statistical and computational performances of BRAIN and the other methods in two typical settings: a low-dimensional setting with $(n,p,p_{a})=(30,2,10^{-3})$ and a high-dimensional setting with $(n,p,p_{a})=(500,250,10^{-4})$.
+ARSRR is omitted from the high-dimensional setting since it is too slow to be a realistic option in this setting.
+The complete results (including ARSRR in moderately challenging settings) are provided in the Appendix D.
+From the results, we can see that BRAIN is significantly faster than ARSRR and PSRR in all settings.
+BRAIN merely requires seconds to generate 1000 acceptable assignments, even in cases where ARSRR or PSRR requires an hour.
+In addition, the computation time of BRAIN does not scale with the sample size $n$, covariate size $p$, or the threshold $p_{a}$ significantly as the other methods, making it an ideal algorithm to use, especially in high-dimensional settings.
 For all methods, the finite-sample biases are negligible, Fisher randomization test controls the type-1 error, and the confidence intervals are valid.
 Compared to CR, all the rerandomization methods reduce the variance of the difference-in-means estimator, increase the statistical power, produce tighter confidence intervals, and have comparable performances in terms of these aspects.
-These results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and substantially reduces the computational cost in simple randomized experiments.
-Table 1: Statistical and computational performance of different methods in simple randomized experiments.
-$n$ $p$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
-Bias SD Size Power Coverage Length
-30 2 CR 6.2 74 10.5 14 89.5 246 $<0.1$
-ARSRR 0.1 51 8.7 19 91.3 177 12.7
-PSRR 1.5 51 10.1 18 89.9 177 50.3
-VNSRR 1.6 52 11.0 17 89.0 177 1.4
-500 250 CR 0.8 201 10.4 73 89.6 660 $<0.1$
-ARSRR 4.0 186 9.2 77 90.8 619 4137.4
-PSRR 4.5 189 9.8 78 90.2 619 275.3
-VNSRR 7.6 199 12.6 73 87.4 615 1.2
+BRAIN generates assignments that are as random as those generated by ARSRR, allowing us to conduct reliable statistical inference (see Appendix D).
+These results demonstrate that BRAIN maintains the appealing statistical properties of rerandomization and substantially reduces the computational cost in simple randomized experiments.
 5.2 Structured Randomized Experiments
 This section considers the three types of structured experiments described in Section 4.
-We simulate the covariates and potential outcomes in the same way as in Section 5.1, except that we set $Y_{i}(1)=Y_{i}(0)+0.15\sqrt{Var[Y_{i}(0)]}$ under the alternative to ensure the power of different methods remains within a moderate range.
+We simulate the covariates and potential outcomes in the same way as in Section 5.1, except that we set $Y_{i}(1)=Y_{i}(0)+0.15\sqrt{\textnormal{Var}[Y_{i}(0)]}$ under the alternative to ensure the power of different methods remains within a moderate range.
 For sequential experiments, we set the number of groups as $K=2$ and consider two group sizes, $n_{k}\in\{100,500\}$.
-We use the strategy proposed by Zhou et al.
-(2018) to determine $a_{k}$.
+We use the strategy proposed by [45] to determine $a_{k}$.
 Specifically, we set $(s_{1},s_{2})=(239,761)$ when $p=50$ and $(s_{1},s_{2})=(264,736)$ when $p=250$, where $a_{k}=\frac{n_{k}}{n_{[k]}}q_{k}$ and $q_{k}$ is the $\frac{1}{s_{k}}$ quantile of a non-central chi-square distribution with $p$ degrees of freedom and a non-central parameter $\frac{n_{[k-1]}}{n_{k}}M_{[k-1]}$.
@@
 For cluster randomized experiments, we set the cluster size as $n_{k}=2$ and consider two options for the number of clusters, $K\in\{100,500\}$.
-We set $p_{a}=P(\chi_{p}^{2}<a)=10^{-3}$ for both stratified randomized experiments and cluster randomized experiments and set $p=n/4$ for all settings in this section.
+We set $p_{a}=P(\chi_{p}^{2}\leq a)=10^{-3}$ for both stratified randomized experiments and cluster randomized experiments and set $p=n/4$ for all settings in this section.
 Table 2 shows the results when $n=1000$.
-The complete results are provided in the Supplement.
+The complete results are provided in Appendix D.
 For clarity, we use a unified name to refer to different method variants.
@@
 The overall findings are consistent with the results in Section 5.1.
-VNSRR substantially reduces the computational burden of rerandomization, especially for high-dimensional datasets and in stratified randomized experiments and cluster randomized experiments, where ARSRR is the only method available to perform rerandomization.
-VNSRR maintains the statistical properties of rerandomization in all the settings and significantly improves statistical inference compared to CR.
-Table 2: Statistical and computational performance of different methods in structured randomized experiments.
+BRAIN substantially reduces the computational burden of rerandomization, especially for high-dimensional datasets and in stratified randomized experiments and cluster randomized experiments, where ARSRR is the only method available to perform rerandomization.
+BRAIN maintains the statistical properties of rerandomization in all the settings and significantly improves statistical inference compared to CR.
+Table 2: Statistical and computational performance of different methods in structured randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
 Scheme Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
-Bias SD Size Power Coverage Length
+Bias SD Size Power CP Len
 Sequential CR 3.2 146 10.7 76 89.3 466 $<0.1$
@@
 PSRR 3.3 135 11.5 82 88.5 429 339.1
-VNSRR 5.8 138 11.4 81 88.6 428 5.8
+BRAIN 5.8 138 11.4 81 88.6 428 5.8
 Stratified CR 3.0 140 10.0 76 90.0 466 $<0.1$
 ARSRR 2.7 131 9.1 82 90.9 434 3386.8
-VNSRR 0.0 130 10.5 83 89.5 436 3.0
+BRAIN 0.0 130 10.5 83 89.5 436 3.0
 Cluster CR 0.9 140 9.8 77 90.2 465 $<0.1$
 ARSRR 6.2 132 9.3 80 90.7 435 3116.2
-VNSRR 3.3 131 10.1 82 89.9 436 3.9
-5.3 A Clinical Trail Example
-In addition to the simulation studies, we analyze a real dataset from a phase 1 clinical trial to illustrate the advantages of VNSRR.
-This dataset was analyzed by Zhu and Liu (2022) as well.
+BRAIN 3.3 131 10.1 82 89.9 436 3.9
+6 A Clinical Trial Example
+In addition to the simulation studies, we analyze a real dataset from a phase 1 clinical trial [2] to illustrate the advantages of BRAIN.
 The goal of this trial was to evaluate the interactions between intravenous methamphetamine and oral reserpine.
 In this trial, 20 participants were randomly assigned to the treatment group, and 10 participants were assigned to the control group.
-We aim to balance better three covariates, age, weight, and pre-treatment heart rate, which are significantly unbalanced in the original study, to better infer the treatment effects on post-treatment heart rate.
+We aim to balance three covariates: age, weight, and pre-treatment heart rate, which are significantly unbalanced in the original study, to better infer the treatment effects on post-treatment heart rate.
 We set $a$ as the 0.001 quantile of $\chi_{3}^{2}$ and generate 10,000 acceptable assignments using each method.
-RR takes approximately 2.3 minutes to complete the task, and PSRR takes more than half an hour.
-In contrast, VNSRR requires only 12 seconds, which demonstrates that VNSRR accelerates the rerandomization process significantly.
-Based on the original dataset, we simulate semi-synthetic data and examine the statistical performance of each method.
-The results are provided in the Supplement.
+ARSRR takes approximately 2.3 minutes to complete the task, and PSRR takes more than half an hour.
+In contrast, BRAIN requires only 12 seconds, which demonstrates that BRAIN accelerates the rerandomization process significantly.
+Based on the original dataset, we simulate semi-synthetic data and examine the statistical performance of each method in Appendix D.
+The results are consistent with our simulation studies, where BRAIN preserves the desired statistical properties of rerandomization.
 7 Conclusion
-In this work, we propose a novel VNSRR method to perform rerandomization efficiently in various types of randomized experiments.
-We derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under VNSRR.
-Our empirical results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and significantly reduces the computational cost compared to existing methods.
-A future direction is to extend the proposed method to experiments with tiers of covariates.
-Multi-objective optimization techniques are required to search for assignments that simultaneously balance the covariates within each tier.
+In this work, we propose a novel BRAIN method to perform rerandomization efficiently in various types of randomized experiments.
+We derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under BRAIN.
+Our empirical results demonstrate that BRAIN maintains the appealing statistical properties of rerandomization and significantly reduces the computational cost compared to existing methods.
+Limitations This work does not address randomized experiments with tiers of covariates.
+Rerandomization for such experiments requires searching for assignments that simultaneously balance the covariates within each tier.
+In future work, we plan to extend our method to this setting using multi-objective optimization techniques.
+Another limitation is that it is unclear whether other integer programming techniques (e.g., branch and bound) can be adapted to develop better rerandomization algorithms.
+We leave this as our future research direction.
 Appendix
-In Appendix A, we provide the pseudocode of the three VNSRR variants.
+In Appendix A, we provide the pseudocode of the three BRAIN variants.
 In Appendix B, we state and prove all the theoretical results.
@@
 In Appendix D, we show the simulation results not included in the main text.
-Appendix A Algorithmic Description of VNSRR Variants
-A.1 Algorithmic Description of SeqVNSRR
-The pseudocode of SeqVNSRR is presented in Algorithm 2.
-As mentioned in the main text, SeqVNSRR performs VNSRR in each experiment stage sequentially while fixing the assignments obtained in previous stages.
-A.2 Algorithmic Description of StratVNSRR
-The pseudocode of StratVNSRR is presented in Algorithm 3.
-StratVNSRR differs from VNSRR in that there are linear constraints within each stratum, $\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk}$.
-Therefore, StratVNSRR only exchanges units in the same stratum to ensure that the aforementioned linear constraints are always satisfied.
-StratVNSRR does not loop over the strata sequentially.
-Instead, after selecting the pairs of units to examine from each stratum, StratVNSRR aggregates the lists of pairs and randomly permutes the combined list.
+Appendix A Algorithmic Description of BRAIN Variants
+A.1 Algorithmic Description of SeqBRAIN
+The pseudocode of SeqBRAIN is presented in Algorithm 2.
+As mentioned in the main text, SeqBRAIN performs BRAIN in each experiment stage sequentially while fixing the assignments obtained in previous stages.
+A.2 Algorithmic Description of StratBRAIN
+The pseudocode of StratBRAIN is presented in Algorithm 3.
+StratBRAIN differs from BRAIN in that there are linear constraints within each stratum, $\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk}$.
+Therefore, StratBRAIN only exchanges units in the same stratum to ensure that the aforementioned linear constraints are always satisfied.
+StratBRAIN does not loop over the strata sequentially.
+Instead, after selecting the pairs of units to examine from each stratum, StratBRAIN aggregates the lists of pairs and randomly permutes the combined list.
 In this way, we ensure that all strata are equally treated, allowing for flexible exploration of the neighborhood without a predetermined ordering of the strata or the units.
-A.3 Algorithmic Description of ClustVNSRR
-The pseudocode of ClustVNSRR is presented in Algorithm 4.
-ClustVNSRR treats each cluster as a whole and uses the clusters as the operational units in the search process.
-In other words, ClustVNSRR essentially conducts VNSRR at the cluster level rather than the unit level.
+A.3 Algorithmic Description of ClustBRAIN
+The pseudocode of ClustBRAIN is presented in Algorithm 4.
+ClustBRAIN treats each cluster as a whole and uses the clusters as the operational units in the search process.
+In other words, ClustBRAIN essentially conducts BRAIN at the cluster level rather than the unit level.
 Recall that we denote the assignment status of the clusters as a $K$-dimensional vector $U$ and all units’ assignment status as a $n$-dimensional vector $W$.
@@
                          Set $M_{[k]}^{(t)}=M^{*}$;
-                         if $M_{[k]}^{(t)}<a_{k}$ then break;
+                         if $M_{[k]}^{(t)}\leq a_{k}$ then break;
+                   end if
+             end for
             if $\widetilde{S}>0$ then
@@
                          Set $t=t+1$;
+                   end for
+             end if
+       end while
+end for
 Output: $W=W^{(t)}$.
-Algorithm 2 Sequential Variable neighborhood searching rerandomization
-Input: Covariates $X$, index sets of units in each stratum $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of units to assign to the two arms in each stratum $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, threshold $a$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$, $L_{total}=\sum_{k}L_{k}$, $S_{total}=\sum_{k}S_{k}$.
+Algorithm 2 Sequential BRAIN
+Input: Covariates $X$, index sets of units in each stratum $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of units to assign to the two arms in each stratum $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, threshold $a$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$, $L_{\textnormal{total}}=\sum_{k}L_{k}$, $S_{\textnormal{total}}=\sum_{k}S_{k}$.
 Set $t=0$;
@@
              Randomly select $L_{k}$ positions of the 0’s in $W^{(t)}$ that belong to $\mathcal{I}_{k}$, denoted by $Z_{k1},\cdots,Z_{kL_{k}}$;
-      Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{L_{total}})$ with random ordering;
-       Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{L_{total}})$.
-The ordering of $(Z_{1},\ldots,Z_{L_{total}})$ is consistent with that of $(O_{1},\ldots,O_{L_{total}})$;
-       Set $\widetilde{S}=S_{total}$;
-       for $l=1,\ldots,L_{total}$ do
+       end for
+      Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{L_{\textnormal{total}}})$ with random ordering;
+       Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{L_{\textnormal{total}}})$.
+The ordering of $(Z_{1},\ldots,Z_{L_{\textnormal{total}}})$ is consistent with that of $(O_{1},\ldots,O_{L_{\textnormal{{total}}}})$;
+       Set $\widetilde{S}=S_{\textnormal{total}}$;
+       for $l=1,\ldots,L_{\textnormal{total}}$ do
              Obtain $W^{*}$ by switching the 1 at position $O_{l}$ and the 0 at position $Z_{l}$ in $W^{(t)}$;
@@
                    Set $M^{(t)}=M^{*}$;
-                   if $M^{(t)}<a$ then break;
+                   if $M^{(t)}\leq a$ then break;
+             end if
+       end for
       if $\widetilde{S}>0$ then
@@
                    Randomly select $S_{k}$ positions of the 0’s in $W^{(t)}$ that are in $\mathcal{I}_{k}$, denoted by $Z_{k1},\cdots,Z_{kS_{k}}$;
-            Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{S_{total}})$;
-             Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{S_{total}})$;
-             for $s=1,\ldots,S_{total}$ do
+             end for
+            Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{S_{\textnormal{total}}})$;
+             Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{S_{\textnormal{total}}})$;
+             for $s=1,\ldots,S_{\textnormal{total}}$ do
                    Obtain $W^{(t+1)}$ by switching the 1 at position $O_{s}$ and the 0 at position $Z_{s}$ in $W^{(t)}$;
@@
                    Set $t=t+1$;
+             end for
+       end if
+end while
 Output: $W=W^{(t)}$.
-Algorithm 3 Stratified Variable neighborhood searching rerandomization
+Algorithm 3 StratBRAIN
 Input: Covariates $X$, index sets of units in each cluster $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of clusters in the treatment arm and the control arm $(K_{t},K_{c})$, threshold $a$, hyperparameter $(L,S)$.
@@
                    Set $U^{(t)}=U^{*}$;
-                   if $M^{(t)}<a$ then break;
+                   if $M^{(t)}\leq a$ then break;
+             end if
+       end for
       if $\widetilde{S}>0$ then
@@
                    Set $t=t+1$;
+             end for
+       end if
+end while
 Output: $W=W(U^{(t)})$.
-Algorithm 4 Cluster Variable neighborhood searching rerandomization
+Algorithm 4 ClustBRAIN
 Appendix B Additional Theoretical Results and Proofs
-VNSRR, StratVNSRR, and ClustVNSRR all assign treatment in an offline manner.
-In contrast, SeqVNSRR performs online allocation and needs to be handled differently.
-We will first establish the theoretical properties of VNSRR, StratVNSRR, and ClustVNSRR, and then obtain the results of SeqVNSRR.
-For the convenience of readers, we first restate the conditions we propose in the main text, which are related to the offline methods.
-Condition 1
+BRAIN, StratBRAIN, and ClustBRAIN all assign treatment in an offline manner.
+In contrast, SeqBRAIN performs online allocation and needs to be handled differently.
+We will first establish the theoretical properties of BRAIN, StratBRAIN, and ClustBRAIN, and then obtain the results of SeqBRAIN.
+We have proposed the following conditions in the main text, which are related to the offline methods.
+Condition 1.
 For $w_{i}=0,1$, $Y_{i}(w_{i})=\beta_{0}+\beta^{T}X_{i}+\tau w_{i}+e_{i}$, where $\beta_{0}+\beta^{T}X_{i}$ is the linear projection of $Y_{i}(0)$ onto $(1,X)$ and $e_{i}$ is the deviation from the linear projection.
-Condition 2
+Condition 2.
 $\widehat{\tau}$ and $\overline{X}_{t}-\overline{X}_{c}$ are normally distributed.
-Condition 3
+Condition 3.
 $n_{t}=n_{c}=\frac{n}{2}$.
-Condition 4
-$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
-Condition 5
+Similar to Condition 3, we also consider the following equal-group-size conditions for stratified randomized experiments and cluster randomized experiments, respectively.
+Condition 4.
+$n_{t_{k}}=n_{c_{k}}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
+Condition 5.
 $K_{t}=K_{c}=\frac{K}{2}$.
-Property 1
-$\mathcal{F}=1-\mathcal{F}$.
-Property 2
-For any $w\in\mathcal{F}$,
-$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
-Property 3
-There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
-$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
-In this section, we first prove Theorem 1 and Theorem 2 based on the abovementioned conditions and properties.
-Theorem 1
-If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
-Theorem 2
-Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
-$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
-where $R^{2}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$.
-We then verify that VNSRR, StratVNSRR, and ClustVNSRR satisfy the requirements by proving the following propositions.
-Their unbiasedness and variance reduction are immediately evident as corollaries of Theorem 1 and Theorem 2.
-Proposition 1
-VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
-Proposition 2
-StratVNSRR satisfies Property 2 and Property 3. Under Condition 4, StratVNSRR satisfies Property 1.
-Proposition 3
-ClustVNSRR satisfies Property 2 and Property 3. Under Condition 5, ClustVNSRR satisfies Property 1.
-We also show that the statistical properties of CR, ARSRR, and PSRR can be established in this way as well based on the following propositions.
-Proposition 4
-CR satisfies Property 1, Property 2 and Property 3.
-Proposition 5
-ARSRR satisfies Property 2 and Property 3. Under Condition 3, ARSRR satisfies Property 1.
-Proposition 6
-PSRR satisfies Property 2 and Property 3. Under Condition 3, PSRR satisfies Property 1.
-Consequently, we will establish the statistical properties of SeqVNSRR by extending the conditions and properties we consider in the offline setting to the online setting.
+We point out three crucial properties of a rerandomization algorithm.
+Property 1.
+(symmetric feasible set) $\mathcal{F}=1-\mathcal{F}$.
+Property 2.
+(uniform initialization) For any $w\in\mathcal{F}$,
+$\mathbb{P}\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
+Property 3.
+(updating rule) There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
+$\mathbb{P}\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
+Any rerandomization algorithm with these properties can yield unbiased treatment effect estimators with reduced variances.
+Theorem 3.
+If $W$ is generated by a rerandomization algorithm satisfying Properties 1-3, then $E[\widehat{\tau}]=\tau$.
+Theorem 4.
+Under Conditions 1 and 2, if $W$ is generated by a rerandomization algorithm satisfying Properties 1-3, then
+$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2},$
+where $R^{2}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}\textnormal{V}_{\textnormal{CR}}}$.
+We shall verify that BRAIN, StratBRAIN, and ClustBRAIN satisfy these properties.
+Proposition 1.
+BRAIN satisfies Properties 2 and 3. Under Condition 3, BRAIN satisfies Property 1.
+Proposition 2.
+StratBRAIN satisfies Properties 2 and 3. Under Condition 4, StratBRAIN satisfies Property 1.
+Proposition 3.
+ClustBRAIN satisfies Properties 2 and 3. Under Condition 5, ClustBRAIN satisfies Property 1.
+Consequently, their unbiasedness and variance reduction results (including Theorems 1 and 2) are immediately evident as corollaries of Theorems 3 and 4.
+In fact, the statistical guarantees of CR, ARSRR, and PSRR can also be established by verifying that they satisfy Properties 1-3.
+We will establish the statistical properties of SeqBRAIN by extending the conditions and properties in the offline setting to the online setting.
 B.1 Proof of Theorem 3
 For any chain $\left(w^{(0)},\ldots,w^{(T)}\right)$ that the rerandomization algorithm may generate, we have
-$P\left(W^{(0)}=w^{(0)}\right)=P\left(W^{(0)}=1-w^{(0)}\right),$
+$\mathbb{P}\left(W^{(0)}=w^{(0)}\right)=\mathbb{P}\left(W^{(0)}=1-w^{(0)}\right),$
 since the initial point $W(0)$ is sampled uniformly from a symmetric space $\mathcal{F}$.
 If we exchange the assignments to the two arms by replacing $W$ with $1-W$, the covariate imbalance between the two arms should remain unchanged. In other words, $M(W)=M(1-W)$. As a result,
-$\hskip 15.0ptP\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)$
+$\hskip 15.0pt\mathbb{P}\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)$
 $=u_{t}\left(M(w^{(0)}),\ldots,M(w^{(t)})\right)$
 $=u_{t}\left(M(1-w^{(0)}),\ldots,M(1-w^{(t)})\right)$
-$=P\left(W^{(t)}=1-w^{(t)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(t-1)}=1-w^{(t-1)}\right),$
+$=\mathbb{P}\left(W^{(t)}=1-w^{(t)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(t-1)}=1-w^{(t-1)}\right),$
 Therefore, for any $w$, we have
-$\hskip 15.0ptP\left(W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)},W^{(T)}=w^{(T)}\right)$
-$=P\left(W^{(0)}=w^{(0)}\right)P\left(W^{(1)}=w^{(1)}\mid W^{(0)}=w^{(0)}\right)\cdots P\left(W^{(T)}=w^{(T)}\mid W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)}\right)$
-$=P\left(W^{(0)}=1-w^{(0)}\right)P\left(W^{(1)}=1-w^{(1)}\mid W^{(0)}=1-w^{(0)}\right)\cdots$
-$\hskip 15.0ptP\left(W^{(T)}=1-w^{(T)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)}\right)$
-$=P\left(W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)},W^{(T)}=1-w^{(T)}\right),$
-We then sum over all possible $\left(w^{(0)},\ldots,w^{(T-1)}\right)$ and obtain $P(W=w^{(T)})=P(W=1-w^{(T)})$. As a result, $P\left(W_{i}=1\right)=P\left(W_{i}=0\right)=1/2$. And we have
-$E[\widehat{\tau}(W)]$ $=E\left[\frac{2}{n}\sum_{i=1}^{n}W_{i}Y_{i}(1)-\frac{2}{n}\sum_{i=1}^{n}\left(1-W_{i}\right)Y_{i}(0)\right]$
-$=\frac{2}{n}\sum_{i=1}^{n}Y_{i}(1)P\left(W_{i}=1\right)-\frac{2}{n}\sum_{i=1}^{n}Y_{i}(0)P\left(W_{i}=0\right)$
+$\hskip 15.0pt\mathbb{P}\left(W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)},W^{(T)}=w^{(T)}\right)$
+$=\mathbb{P}\left(W^{(0)}=w^{(0)}\right)\mathbb{P}\left(W^{(1)}=w^{(1)}\mid W^{(0)}=w^{(0)}\right)\cdots\mathbb{P}\left(W^{(T)}=w^{(T)}\mid W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)}\right)$
+$=\mathbb{P}\left(W^{(0)}=1-w^{(0)}\right)\mathbb{P}\left(W^{(1)}=1-w^{(1)}\mid W^{(0)}=1-w^{(0)}\right)\cdots$
+$\hskip 15.0pt\mathbb{P}\left(W^{(T)}=1-w^{(T)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)}\right)$
+$=\mathbb{P}\left(W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)},W^{(T)}=1-w^{(T)}\right),$
+We then sum over all possible $\left(w^{(0)},\ldots,w^{(T-1)}\right)$ and obtain $\mathbb{P}(W=w^{(T)})=\mathbb{P}(W=1-w^{(T)})$.
+As a result, $\mathbb{P}\left(W_{i}=1\right)=\mathbb{P}\left(W_{i}=0\right)=1/2$.
+And we have
+$\mathbb{E}[\widehat{\tau}(W)]$ $=\mathbb{E}\left[\frac{2}{n}\sum_{i=1}^{n}W_{i}Y_{i}(1)-\frac{2}{n}\sum_{i=1}^{n}\left(1-W_{i}\right)Y_{i}(0)\right]$
+$=\frac{2}{n}\sum_{i=1}^{n}Y_{i}(1)\mathbb{P}\left(W_{i}=1\right)-\frac{2}{n}\sum_{i=1}^{n}Y_{i}(0)\mathbb{P}\left(W_{i}=0\right)$
 $=\tau.$
@@
 Since the rerandomization algorithm only affects $\overline{X}_{t}-\overline{X}_{c}$ and does not affect $\overline{e}_{t}-\overline{e}_{c}$, we have
-$Var[\widehat{\tau}]$ $=Cov\left[\beta^{T}(\overline{X}_{t}-\overline{X}_{c})\beta\right]+Var\left[\overline{e}_{t}-\overline{e}_{c}\right]$
-$=\beta^{T}Cov\left[\overline{X}_{t}-\overline{X}_{c}\right]\beta+\left(1-R^{2}\right)V_{CR}$
+$\textnormal{Var}[\widehat{\tau}]$ $=\textnormal{Cov}\left[\beta^{T}(\overline{X}_{t}-\overline{X}_{c})\beta\right]+\textnormal{Var}\left[\overline{e}_{t}-\overline{e}_{c}\right]$
+$=\beta^{T}\textnormal{Cov}\left[\overline{X}_{t}-\overline{X}_{c}\right]\beta+\left(1-R^{2}\right)\textnormal{V}_{\textnormal{CR}}$
 We then bound the first term in the variance of $\widehat{\tau}$.
 Recall that
-$M=\frac{n}{n_{t}n_{c}}(W-\frac{n_{t}}{n}1_{n})^{T}XS_{XX}^{-1}X^{T}(W-\frac{n_{t}}{n}1_{n}).$
+$M=\frac{n}{n_{t}n_{c}}\left(W-\frac{n_{t}}{n}1_{n}\right)^{T}XS_{XX}^{-1}X^{T}\left(W-\frac{n_{t}}{n}1_{n}\right).$
 We define
-$Z=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}(W-\frac{n_{t}}{n}1_{n})=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}(\overline{X}_{t}-\overline{X}_{c}).$
+$Z=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}X^{T}\left(W-\frac{n_{t}}{n}1_{n}\right)=\sqrt{\frac{n}{n_{t}n_{c}}}S_{XX}^{-1/2}(\overline{X}_{t}-\overline{X}_{c}).$
 Then we have $M=Z^{T}Z$.
@@
 We note that the rows of $S_{XX}^{-1/2}X^{T}$ are exchangeable since the covariates are normalized. As a result, $\{Z_{j}\}_{1\leq j\leq p}$ are exchangeable, and we have
-$Var[Z_{j}]=E[Z_{j}^{2}]=\frac{1}{p}E\left[\sum_{j=1}^{p}Z_{j}^{2}\right]=\frac{1}{p}E[M]\leq\frac{a}{p}.$
+$\textnormal{Var}[Z_{j}]=E[Z_{j}^{2}]=\frac{1}{p}E\left[\sum_{j=1}^{p}Z_{j}^{2}\right]=\frac{1}{p}E[M]\leq\frac{a}{p}.$
 If we flip the sign of the $j$-th normalized covariate, the sign of $Z_{j}$ is also flipped, but the mapping $M(W)$ remains unchanged.
@@
 And we have
-$Cov[Z_{j},Z_{l}]$ $=E[Z_{j}Z_{l}]$
+$\textnormal{Cov}[Z_{j},Z_{l}]$ $=E[Z_{j}Z_{l}]$
 $=E[E\left[Z_{j}Z_{l}\mid Z_{l}\right]]$
@@
 $=0.$
-Therefore, $Cov[Z]=Var[Z_{j}]I_{p}$, and
-$\beta^{T}Cov[\overline{X}_{t}-\overline{X}_{c}]\beta$ $=\beta^{T}Cov\left[\sqrt{\frac{n_{t}n_{c}}{n}}S_{XX}^{1/2}Z\right]\beta$
-$=\frac{n_{t}n_{c}}{n}\beta^{T}S_{XX}^{1/2}Cov[Z]S_{XX}^{1/2}\beta$
-$=\frac{n_{t}n_{c}}{n}Var[Z_{j}]\beta^{T}S_{XX}\beta$
-$=Var[Z_{j}]R^{2}V_{CR}$
-$\leq\frac{a}{p}R^{2}V_{CR}.$
+Therefore, $\textnormal{Cov}[Z]=\textnormal{Var}[Z_{j}]I_{p}$, and
+$\beta^{T}\textnormal{Cov}[\overline{X}_{t}-\overline{X}_{c}]\beta$ $=\beta^{T}\textnormal{Cov}\left[\sqrt{\frac{n_{t}n_{c}}{n}}S_{XX}^{1/2}Z\right]\beta$
+$=\frac{n_{t}n_{c}}{n}\beta^{T}S_{XX}^{1/2}\textnormal{Cov}[Z]S_{XX}^{1/2}\beta$
+$=\frac{n_{t}n_{c}}{n}\textnormal{Var}[Z_{j}]\beta^{T}S_{XX}\beta$
+$=\textnormal{Var}[Z_{j}]R^{2}\textnormal{V}_{\textnormal{CR}}$
+$\leq\frac{a}{p}R^{2}\textnormal{V}_{\textnormal{CR}}.$
 Consequently,
-$Var[\widehat{\tau}]\leq\frac{a}{p}R^{2}V_{CR}+(1-R^{2})V_{CR}.$
+$\textnormal{Var}[\widehat{\tau}]\leq\frac{a}{p}R^{2}\textnormal{V}_{\textnormal{CR}}+(1-R^{2})\textnormal{V}_{\textnormal{CR}}.$
 Rearranging the terms, we obtain
-$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
+$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
 B.3 Proof of the Propositions
@@
 $\sum_{i}(1-W_{i})=n-n_{t}=n_{c}=n_{t}.$
-This implies that $1-W\in\mathcal{F}_{0}$ as well. Thus, $\mathcal{F}_{0}=1-\mathcal{F}_{0}$. Since CR, ARSRR, PSRR, and VNSRR treat $\mathcal{F}_{0}$ as the feasible set, they all satisfy Property 1.
-StratVNSRR and ClustVNSRR conduct rerandomization with more constraints imposed by the structure of the experiments. We show that these constraints are also symmetric in nature.
-We denote the feasible set of StratVNSRR and ClustVNSRR as $\mathcal{F}_{Strat}$ and $\mathcal{F}_{Clust}$, respectively.
+This implies that $1-W\in\mathcal{F}_{0}$ as well. Thus, $\mathcal{F}_{0}=1-\mathcal{F}_{0}$. Since CR, ARSRR, PSRR, and BRAIN treat $\mathcal{F}_{0}$ as the feasible set, they all satisfy Property 1.
+StratBRAIN and ClustBRAIN conduct rerandomization with more constraints imposed by the structure of the experiments. We show that these constraints are also symmetric in nature.
+We denote the feasible set of StratBRAIN and ClustBRAIN as $\mathcal{F}_{Strat}$ and $\mathcal{F}_{Clust}$, respectively.
 Under Condition 4, for any $W\in\mathcal{F}_{Strat}$,
@@
 This implies that $1-W\in\mathcal{F}_{Clust}$ and $\mathcal{F}_{Clust}=1-\mathcal{F}_{Clust}$.
-Therefore, both StratVNSRR and ClustVNSRR satisfy Property 1.
+Therefore, both StratBRAIN and ClustBRAIN satisfy Property 1.
 All of the methods satisfy Property 2 since they all start with a random assignment by uniformly sampling from the feasible set.
@@
 Therefore, both ARSRR and PSRR satisfy Property 3.
-For VNSRR and its variants, $(M(w^{(0)}),\ldots,M(w^{(t-1)}))$ determines whether $W^{(t)}$ is obtained through the local search step or the shaking step.
+For BRAIN and its variants, $(M(w^{(0)}),\ldots,M(w^{(t-1)}))$ determines whether $W^{(t)}$ is obtained through the local search step or the perturbation step.
 In the local search step, we have
 $u_{t}=1\{M(w^{(t)})<M(w^{(t-1)})\}.$
-And in the shaking step, $u_{t}$ is a constant which depends on the number of adjacent feasible solutions.
-Therefore, VNSRR satisfies Property 3.
-B.4 Theoretical Results for SeqVNSRR
-In sequential experiments, we have the following Condition 4, which corresponds to the Condition 3 in the non-sequential settings.
-Condition 4
+And in the perturbation step, $u_{t}$ is a constant which depends on the number of adjacent feasible solutions.
+Therefore, BRAIN satisfies Property 3.
+B.4 Theoretical Results for SeqBRAIN
+We consider the following equal-group-size condition for sequentially randomized experiments.
+We note that it has the same form as the condition for stratified randomized experiments due to the similarity between the notations in the two settings.
+Condition 4.
 $n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
-We also have the corresponding version of Property 1-3, stated as follows.
-Property 4
+We also have the corresponding version of Properties 1-3, stated as follows.
+Property 4.
 For any $k$, $\mathcal{F}_{[k]}=1-\mathcal{F}_{[k]}$.
-Property 5
+Property 5.
 For any $k$ and $w\in\mathcal{F}_{[k]}$,
 $P\left(W_{[k]}^{(0)}=w\right)=\frac{1}{|\mathcal{F}_{[k]}|}.$
-Property 6
+Property 6.
 There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $k$, $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain for the $k$-th stage,
@@
 Based on these sufficient conditions, we obtain that
-Theorem 3
-If $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then $E[\widehat{\tau}]=\tau$.
-Theorem 4
-Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 4, Property 5 and Property 6, then
-$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a_{K}}{p}\right)R^{2},$
-We omit the proofs of Theorem 3 and Theorem 4, since they are almost the same as the proofs of Theorem 1 and Theorem 2, except that the chain $(W^{(0)},\ldots,W^{(T)})$ needs to be decomposed into $K$ parts corresponding to the $K$ stages.
+Theorem 5.
+If $W$ is generated by a rerandomization algorithm satisfying Properties 4-6, then $E[\widehat{\tau}]=\tau$.
+Theorem 6.
+Under Conditions 1 and 2, if $W$ is generated by a rerandomization algorithm satisfying Properties 4-6, then
+$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a_{K}}{p}\right)R^{2},$
+We omit the proofs of Theorems 5 and 6, since they are similar to the proofs of Theorems 1 and 2, except that the chain $(W^{(0)},\ldots,W^{(T)})$ needs to be decomposed into $K$ parts corresponding to the $K$ stages.
 Each part can be analyzed in the same way as in simple randomized experiments.
-We have verified that VNSRR satisfies Property 1-3.
-We note that SeqVNSRR can be viewed as a special version of VNSRR in each of its stages.
-As a result, SeqVNSRR satisfies the sequential version of Property 1-3.
-Proposition 7
-SeqVNSRR satisfies Property 5 and Property 6. Under Condition 4, SeqVNSRR satisfies Property 4.
-The unbiasedness and variance reduction properties of SeqVNSRR are then immediately evident as corollaries of Theorem 3 and Theorem 4.
+We have verified that BRAIN satisfies Properties 1-3.
+We note that SeqBRAIN can be viewed as a special version of BRAIN in each of its stages.
+As a result, SeqBRAIN satisfies the sequential version of Properties 1-3.
+Proposition 4.
+SeqBRAIN satisfies Properties 5 and 6. Under Condition 4, SeqBRAIN satisfies Property 4.
+The unbiasedness and variance reduction properties of SeqBRAIN are then immediately evident as corollaries of Theorems 5 and 6.
 Appendix C Randomization-based Inference
@@
 The p-value is then calculated by comparing $\{\widehat{\tau}(W^{b})\}_{1\leq b\leq B}$ with the observed treatment effect estimator $\widehat{\tau}(W^{obs})$, $pv=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b})|\geq|\widehat{\tau}(W^{obs})|\}$
-As pointed out by Zhu and Liu (2022), randomization-based confidence intervals can be constructed by inverting Fisher randomization tests.
+As pointed out by [46], randomization-based confidence intervals can be constructed by inverting Fisher randomization tests.
 To construct a lower confidence bound for $\tau$, we consider testing $H_{0}:Y_{i}(1)-Y_{i}(0)=\theta,i=1,\ldots,n$, versus $H_{1}:Y_{i}(1)-Y_{i}(0)>\theta,i=1,\ldots,n$.
@@
 The p-value is then calculated as $pv(\theta)=\frac{1}{B}\sum_{b=1}^{B}1\{|\widehat{\tau}(W^{b},\theta)|\geq|\widehat{\tau}(W^{obs})|\}$.
-Luo et al.
-(2021) proved that $\theta_{l}=\sup\{\theta:pv(\theta)\leq\alpha\}$ is a lower confidence bound for $\tau$ with confidence level $\alpha$.
-And Zhu and Liu (2022) proposed an efficient approach to solve for $\theta_{l}$, which we adopt in our numerical experiments.
+[25] proved that $\theta_{l}=\sup\{\theta:pv(\theta)\leq\alpha\}$ is a lower confidence bound for $\tau$ with confidence level $\alpha$.
+And [46] proposed an efficient approach to solve for $\theta_{l}$, which we adopt in our numerical experiments.
 Similarly, we can construct an upper confidence bound $\theta_{u}$ with confidence level $\alpha$.
@@
 Appendix D Simulation Results
-In Section D.1, we provide the complete results of the simulation studies in the main text.
-In Section D.2, we examine the randomness of assignments generated by each method.
-In Section D.3, we generate semi-synthetic data based on the clinical trial dataset and compare the statistical performance of each method.
+In D.1, we provide the complete results of the simulation studies in the main text.
+In D.2, we examine the randomness of assignments generated by each method.
+In D.3, we compare different combinations of hyperparameters.
+In D.4, we generate semi-synthetic data based on the clinical trial dataset and compare the statistical performance of each method.
 D.1 Statistical and Computational Performances
@@
 The overall findings are similar to the conclusion in the main text.
-VNSRR substantially reduces the computational burden of rerandomization and maintains the appealing statistical properties of rerandomization.
+BRAIN substantially reduces the computational burden of rerandomization and maintains the appealing statistical properties of rerandomization.
 D.2 Randomness Metric
-Similar to Zhu and Liu (2022), we denote $L_{n}$ as the largest eigenvalue of the covariance matrix of $12W-$ 1 and use $L_{n}$ to measure the randomness of assignments generated by each method.
-As pointed out by Zhu and Liu (2022), larger $L_{n}$ implies less randomness.
+Similar to [46], we denote $L_{n}$ as the largest eigenvalue of the covariance matrix of $12W-$ 1 and use $L_{n}$ to measure the randomness of assignments generated by each method.
+As pointed out by [46], larger $L_{n}$ implies less randomness.
 Table 5 and Table 6 show the results in simple experiments and structured experiments, respectively.
@@
 The three rerandomization methods perform similarly in all settings.
-Therefore, one can safely perform randomization-based statistical inference based on the assignments generated by VNSRR without the risk of relying on a small set of unique assignments.
-D.3 Semi-synthetic Data
-We mimic the set-ups in Zhu and Liu (2022) and simulate semi-synthetic data based on the clinical trial dataset.
+Therefore, one can safely perform randomization-based statistical inference based on the assignments generated by BRAIN without the risk of relying on a small set of unique assignments.
+D.3 Hyperparameters
+There are $n$ units, hence we can swap at most $n/2$ disjoint pairs in an iteration.
+So, the maximum value of $L$ and $S$ is $n/2$.
+The minimum value of $L$ and $S$ is 1.
+We also consider the maximum value times 0.25, 0.5, and 0.75.
+Overall, there are $5\times 5=25$ combinations in total.
+Figure 1 shows the computational time of BRAIN with different combinations of $L$ and $S$ under the setting where $(n,p,p_{a})=(30,15,10^{-3})$.
+It demonstrates that our choice of $(L,S)$ is among the top tier.
+D.4 Semi-synthetic Data
+We mimic the set-ups in [46] and simulate semi-synthetic data based on the clinical trial dataset.
 Since we only observe half of the potential outcomes in the original dataset, we impute the counterfactual potential outcomes under the sharp null.
-We then compare VNSRR and the existing methods in the same way as in our simulation studies.
+We then compare BRAIN and the existing methods in the same way as in our simulation studies.
 The results are shown in Table 7 and are similar to our findings in the simulation studies.
-VNSRR significantly accelerates rerandomization and has comparable performances in other aspects compared to ARSRR and PSRR.
-Table 3: Statistical and computational performance of different methods in simple randomized experiments.
+BRAIN significantly accelerates rerandomization and has comparable performances in other aspects compared to ARSRR and PSRR.
+Table 3: Statistical and computational performance of different methods in simple randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
 $n$ $p$ $p_{a}$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
-Bias SD Size Power Coverage Length
+Bias SD Size Power CP Len
 30 2 $10^{-3}$ CR 6.2 74 10.5 14 89.5 246 $<0.1$
@@
 PSRR 1.5 51 10.1 18 89.9 177 48.1
-VNSRR 1.6 52 11.0 17 89.0 177 1.4
+BRAIN 1.6 52 11.0 17 89.0 177 1.4
 15 $10^{-3}$ CR 9.2 199 9.3 18 90.7 678 $<0.1$
@@
 PSRR 3.8 156 10.6 22 89.4 547 1.6
-VNSRR 6.8 152 10.4 22 89.6 547 0.3
+BRAIN 6.8 152 10.4 22 89.6 547 0.3
 100 2 $10^{-3}$ CR 1.4 42 11.7 29 88.3 132 $<0.1$
@@
 PSRR 0.3 29 9.6 40 90.4 94 2.1
-VNSRR 0.4 27 8.5 40 91.5 94 0.5
+BRAIN 0.4 27 8.5 40 91.5 94 0.5
 $10^{-4}$ CR 1.4 42 11.7 29 88.3 132 $<0.1$
@@
 PSRR 0.1 28 10.1 40 89.9 94 35.7
-VNSRR 0.3 28 10.2 41 89.8 94 4.2
+BRAIN 0.3 28 10.2 41 89.8 94 4.2
 50 $10^{-3}$ CR 1.0 207 11.3 35 88.7 663 $<0.1$
@@
 PSRR 5.1 170 10.1 42 89.9 574 8.2
-VNSRR 2.1 173 10.1 40 89.9 573 0.2
+BRAIN 2.1 173 10.1 40 89.9 573 0.2
 $10^{-4}$ CR 1.0 207 11.3 35 88.7 663 $<0.1$
@@
 PSRR 2.8 167 9.9 43 90.1 560 27.0
-VNSRR 6.4 163 9.1 42 90.9 558 0.3
+BRAIN 6.4 163 9.1 42 90.9 558 0.3
 500 2 $10^{-3}$ CR 1.0 18 8.7 74 91.3 59 $<0.1$
@@
 PSRR 0.6 12 9.1 94 90.9 42 1.6
-VNSRR 0.1 13 10.5 93 89.5 42 0.5
+BRAIN 0.1 13 10.5 93 89.5 42 0.5
 $10^{-4}$ CR 1.0 18 8.7 74 91.3 59 $<0.1$
@@
 PSRR 0.5 13 11.2 92 88.8 42 6.4
-VNSRR 0.2 12 11.3 92 88.7 42 1.7
+BRAIN 0.2 12 11.3 92 88.7 42 1.7
 250 $10^{-3}$ CR 0.8 201 10.4 73 89.6 660 $<0.1$
@@
 PSRR 4.5 189 9.8 78 90.2 619 275.3
-VNSRR 7.6 199 12.6 73 87.4 615 1.2
+BRAIN 7.6 199 12.6 73 87.4 615 1.2
 $10^{-4}$ CR 0.8 201 10.4 88 89.6 660 $<0.1$
 PSRR 4.3 187 10.3 91 89.7 611 2878.8
-VNSRR 5.7 181 9.8 92 90.2 608 1.3
-Table 4: Statistical and computational performance of different methods in structured randomized experiments.
+BRAIN 5.7 181 9.8 92 90.2 608 1.3
+Table 4: Statistical and computational performance of different methods in structured randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
 Scheme $n$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
-Bias SD Size Power Coverage Length
+Bias SD Size Power CP Len
 Sequential 200 CR 5.5 140 10.3 28 89.7 469 $<0.1$
@@
 PSRR 1.7 118 10.4 36 89.6 391 14.1
-VNSRR 0.9 116 10.2 35 89.8 390 0.9
+BRAIN 0.9 116 10.2 35 89.8 390 0.9
 1000 CR 3.2 146 10.7 76 89.3 466 $<0.1$
@@
 PSRR 3.3 135 11.5 82 88.5 429 339.1
-VNSRR 5.8 138 11.4 81 88.6 428 5.8
+BRAIN 5.8 138 11.4 81 88.6 428 5.8
 Stratified 200 CR 5.2 140 10.3 26 89.7 467 $<0.1$
 ARSRR 4.5 121 10.6 35 89.4 401 148.0
-VNSRR 6.1 124 10.3 32 89.7 402 0.5
+BRAIN 6.1 124 10.3 32 89.7 402 0.5
 1000 CR 3.0 140 10.0 76 90.0 466 $<0.1$
 ARSRR 2.7 131 9.1 82 90.9 434 3386.8
-VNSRR 0.0 130 10.5 83 89.5 436 3.0
+BRAIN 0.0 130 10.5 83 89.5 436 3.0
 Cluster 200 CR 3.0 141 10.9 27 89.1 479 $<0.1$
 ARSRR 3.3 116 8.8 32 91.2 405 126.4
-VNSRR 6.4 116 8.9 30 91.1 405 0.8
+BRAIN 6.4 116 8.9 30 91.1 405 0.8
 1000 CR 0.9 140 9.8 77 90.2 465 $<0.1$
 ARSRR 6.2 132 9.3 80 90.7 435 3116.2
-VNSRR 3.3 131 10.1 82 89.9 436 3.9
+BRAIN 3.3 131 10.1 82 89.9 436 3.9
 Table 5: $L_{n}$ (×$\times 10^{-2}$) of different methods in simple randomized experiments.
 $n$ $p$ $p_{a}$ Method
-CR ARSRR PSRR VNSRR
+CR ARSRR PSRR BRAIN
 30 2 $10^{-3}$ 137 158 161 158
@@
 Scheme $n$ Method
-CR ARSRR PSRR VNSRR
+CR ARSRR PSRR BRAIN
 Sequential 200 208 251 252 253
@@
 1000 576 606 NA 604
-Table 7: Statistical and computational performance of different methods when applied to semi-synthetic data.
+Figure 1: Comparison of different hyperparameters.
+Table 7: Statistical and computational performance of different methods when applied to semi-synthetic data. SD: standard deviation. CP: coverage probability. Len: length.
 Method Statistical Inference (×$\times 10^{-3}$) Computation Time (second) $L_{n}$
-Bias SD Size Coverage Length
+Bias SD Size CP Len
 CR 9.7 440 4.4 95.6 1804 $<0.1$ 122
@@
 PSRR 0.6 319 4.8 95.2 1346 219.9 142
-VNSRR 14.4 336 5.3 94.7 1338 1.2 139
+BRAIN 14.4 336 5.3 94.7 1338 1.2 139